### v1

# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

Abstract.
The aim of this article is twofold. First, we develop a unified framework for viscosity solutions to both first-order Hamilton-Jacobi equations and semilinear Hamilton–Jacobi equations driven by the idiosyncratic operator. Second, we establish a vanishing-viscosity limit—extending beyond the classical control-theoretic setting—for solutions of semilinear Hamilton–Jacobi equations, proving their convergence to the corresponding first-order solution as the idiosyncratic noise vanishes. Our approach provides an optimal convergence rate.
We also present some results of independent interest. These include existence theorems for the first-order equation, obtained through an appropriate Hopf–Lax representation, and a useful description of the action of the idiosyncratic operator on geodesically convex functions.
1991 Mathematics Subject Classification:
49N80, 35R15, 49022
1. Introduction
The starting objective of this work is to treat in a unified approach the following two Hamilton-Jacobi equations (HJE)s defined on the Wasserstein space $\mathcal{P}_{2}(\mathbb{R}^{d})$: the first order HJE
[…]
and the semilinear HJE with idiosyncratic noise
(1.2) $\displaystyle-\partial_{t}U(t,\mu)\;+\;\int_{\mathbb{R}^{d}}H\bigl(x,\partial_{\mu}U(t,\mu,x),\mu\bigr)\,d\mu(x)\;-\;\int_{\mathbb{R}^{d}}\mathrm{tr}\!\left[a(\mu,x)\,\nabla_{x}\partial_{\mu}U(t,\mu,x)\right]d\mu(x)=0,$ $\text{ in }[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$
where, $T>0$, $H:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is the Hamiltonian, and $a:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to M_{d\times d}(\mathbb{R})$ is the diffusion matrix that will be assumed to be positive definite. Usually we will consider the final time condition $U(T,\cdot)=\mathcal{G}(\cdot)$ where $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is the terminal datum. Here, $\partial_{\mu}U(t,\mu,x)$ stands for the Wasserstein gradient of $U$ at the point $(t,\mu,x)$.
In recent years, a viscosity-solution theory has been developed to address HJEs in the space of probability measures, ensuring well-posedness for this class of nonlinear equations. Classically, uniqueness and stability follow from the comparison principle, while the existence theory has been mainly established in the convex setting, identifying the unique viscosity solution with the value function of an optimal control problem on the space of probability measures. This framework yields the rigorous connection between HJEs on $\mathcal{P}_{2}(\mathbb{R}^{d})$ and Mean Field Control problems.
In this paper, we develop an approach that enables us to treat these two equations using a unified definition of a viscosity solution. In particular, we establish an asymptotic relation between the viscosity solutions of (1.1) and (1.2) through a vanishing-viscosity limit. To the best of our knowledge, this is the first work to provide such a perturbative link without imposing regularity assumptions on the functions $U^{\varepsilon}$, the viscosity solutions to (1.2) when $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$, apart from those inherent to the control setting. In [tangpi2022probabilisticapproachvanishingviscosity], a vanishing viscosity result is obtained also in the presence of a common noise operator, in the convex case; however, that analysis relies on additional regularity assumptions on the solutions.
1.1. Overview on the existing literature
The study of HJEs on Wasserstein spaces has gained a lot of interest, in these last years, thanks to their connections with Mean Field Control, Mean Field Games, and the associated Master Equation. Mean Field Games theory has been introduced by Lasry and P.L.Lions [LLMeanFG], and independently by Huang, Caines and Malhamé [MeanFieldCaines]. It has been subsequently developed in Lions’ courses at Collège de France [College] and it has attracted the interest of a large group of researchers. The well-posedness theory for solutions of HJEs in the space of probability measures is one of the objectives of this theory due to its connection with mean-field particle approximation [cardaliaguet2022regularityvaluefunctionquantitative, cecchin2025quantitativeconvergencemeanfield], the control of Mckean-Vlasov equation [CarmonaDelarue, CarmonaDelarueLachapelle2013], and large deviation theory [FengKurtz2006].
In the study of HJEs in $\mathcal{P}_{2}(\mathbb{R}^{d})$, various notions of viscosity solutions have been proposed in the literature (possibly in presence of a common noise) with the shared aim of extending the classical finite-dimensional framework to this more complex setting. However, this new context presents significant challenges. The geometry of $\mathcal{P}_{2}(\mathbb{R}^{d})$ endowed with the Wasserstein distance $W_{2}$, originally exploited in the seminal work by Otto in [Otto], and later developed in the monograph by Ambrosio et al. [AGS], exhibits a singular and non linear structure. As a result, performing differential calculus in $\mathcal{P}_{2}(\mathbb{R}^{d})$ involves the non-trivial task of developing a suitable notion of differentiation in a non-smooth realm. In [AGS], a subdifferential calculus has been introduced, allowing to give a finer description of the Gradient Flows in the space of probability measures. In particular, superdifferentiability of the squared Wasserstein distance is established. We remark that the lack of (local) differentiability of the distance is a common issue in the theory of HJEs in infinite dimensional spaces. The infinite dimensional theory has been firstly studied by Crandall and P.L. Lions in the series of papers [CrandallLions1984, CrandallLions1986a, CrandallLions1986b, CrandallLions1990, CrandallLions1991], where, to treat HJEs in Banach spaces, it is assumed the structural hypothesis of the existence of a smooth metric induced by the assigned norm.
To overcome these difficulties in the Wasserstein setting, several strategies have been explored.
  • Penalization via entropy terms. The singular geometry of the space is reflected by the non-differentiability of the squared Wasserstein distance $W_{2}$ at “non-regular” measures (see [gigc, Section  2]). As described by Daudin and Seeger in [DS], by Bertucci and P.L. Lions in [BertucciApprox], and in the series of papers of the second author et al. [CKTExi],[CKTComp], wellposedness for semilinear HJEs in the space of probability measures can be obtained by adding an entropic penalization that localizes the equation at points of differentiability of the squared distance. The semilinearity assumption is fundamental since it helps to absorb the variational contribution of the entropy penalization in the HJE. These approaches are reminiscent of those introduced by Tataru [Tataru1992, Tataru1994] to treat singular HJBs in Banach spaces, and the one developed by Feng and Katsoulakis [FengKatsoulakis2009] to handle gradient flows in infinite dimensions.
    Extending the theory to general, even degenerate and not variational, parabolic equation is an interesting problem out of the scope of this work.
  • Regularization via weaker metrics. The lack of regularity of the distance can be addressed by replacing it with a more regular alternative, following a program analogous to that developed in Banach spaces. One possibility is to modify the geometric structure of the space by introducing weaker yet sufficiently smooth metrics, as discussed by Soner and Yan in [STorus]. A similar idea appears in Cosso et al. [cosso2022masterbellmanequationwasserstein], where a gauge function is employed and common noise is also handled. Along the same lines, Bayraktar et al. [bayraktar2023comparisonviscositysolutionsclass] introduce a clever change of variables that allows one to treat common noise as a finite-dimensional operator, to which an Ishii–Lions lemma (see [crandall1992usersguideviscositysolutions]) can be applied. However, this approach has a drawback: the conditions required to guarantee the comparison principle become difficult to satisfy. Indeed, the weaker the metric, the more regularity is demanded of the data. This trade-off is absent in the penalization approach, where the assumptions on the Hamiltonian closely match those in the finite-dimensional setting.
  • Lifting to a Hilbert space. In the courses at the Collège de France [College], P.L. Lions proposed to lift the HJE to a more familiar and smoother space, such as a Hilbert space, where an extensive viscosity solution theory has been developed in the 80’s and 90’s. Under this viewpoint, the Wasserstein space is realized as a (metric) submersion, and can be regarded as the space of laws of squared-integrable random variables. creates a bridge between the intrinsic calculus developed in [AGS] and extrinsic calculus via the so called Lions’ derivative. See also Carmona and Delarue [CarmonaDelarue, Chapter  5], Bertucci [BSOT], and Gangbo and Tudorascu [GANGBO2019119], where the use of Lions’ lift to treat HJE over $\mathcal{P}_{2}(\mathbb{R}^{d})$ has been deeply treated. For HJEs without idiosyncratic noise terms, this lift provides a natural strategy to give a well-posed notion of viscosity solution in the space of probability measures [BSOT]. However, as already described in [DS, Rmk 2.20], this approach does not adapt well to treat idiosyncratic noise, due to the singularity of the lifted idiosyncratic operator.
  • Viewing the Wasserstein space as a Positively Curved (PC) space. Some authors have instead chosen to develop a more general and abstract theory of HJEs in merely (lenght) metric spaces, see [AMBFeng, GangboSwiech2015_metricViscosity]. Althoug this approach is elegant, it restricts the class of admissible Hamiltonians to the ones depending only on the metric slope, therefore not sensible to directional variations of the unknown function. A further step was to observe that $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a PC space, see [AGS, Appendix]. This geometric property allows one to define a notion of tangent bundle endowed with a weak scalar product. A test-function approach can then be employed to obtain well-posedness. See also[JerhaouiAussedatZidani2024] for an application of this construction. As observed in [DS], the link between test functions and the subdifferential approach is not yet understood. For this reason, in the idiosyncratic case we adopt a more classical notion of solution, leaving the investigation of the relationship between the two frameworks to future work.
  • Superdifferentiability of the Wasserstein distance. In [BSOT], the author realizes that a superdifferentiability of the Wasserstein distance is sufficient to establish a comparison principle for (1.1). In particular, he proposes a notion that ensures both the superdifferentiability of the squared Wasserstein distance and existence results based on the Hilbertian lift.
1.2. Our main results and contributions.
We describe here the main contributions that we add to the literature:
  • (1)
    Unifying notion of superdifferential:
    In this work, we provide a careful analysis of the weak Riemannian structure in the Wasserstein space, in order to select a superdifferential $\partial^{+}$ with the following properties:
    • (i)
      The Wasserstein distance squared is superdifferentiable, following the approach in [BSOT].
    • (ii)
      $\partial^{+}$ is weaker than the definition of superdifferential given in [BSOT].
    • (iii)
      $\partial^{+}$ coincides with the superdifferential adopted in [DS, Definition  2.5] when restricted to measures that are absolutely continuous w.r.t. the Lesbesgue measure.
    The appropriate notion of superdifferential $\partial^{+}$ is
    $\partial^{+}=\partial_{S}^{+}\cap\text{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),$
    where $\partial_{S}$ is the the strong differential, see Definition 2.7, and $\text{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the geometric Tangent space at a measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, see Definition 2.2 Indeed, $\partial^{+}$ is a subset of the superdifferential defined in [BSOT]. On the other hand, in $\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$, the set absolutely continuous measures w.r.t. the Lesbesgue measure, the baricentric projection yields
    $(\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),W_{\mu})\simeq_{\text{Pr}}(\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),\|\cdot\|_{L^{2}_{\mu}})=(\overline{\left\{\nabla\phi\colon\phi\in C^{\infty}_{c}(\mathbb{R}^{d})\right\}}^{L^{2}_{\mu}(\mathbb{R}^{d})},\|\cdot\|_{L^{2}_{\mu}(\mathbb{R}^{d})})\quad\forall\mu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d}).$
    This identification allows to relate our notion of super differential with the one introduced in [DS], namely:
    $\partial^{+}_{S}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\simeq_{\mathrm{Pr}}\partial^{+}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}).$
    We also refer to Lemma 2.12 for more details. Although this result is not new in the literature, see [GANGBO2019119, Theorem  3.14], we provide here an alternative proof using the weak Riemannian formalism introduced in [GigliPhd] and adopted in this work.
    Moreover, in Proposition (2.14) we show that Wasserstein distance squared is superdifferentiable w.r.t to $\partial^{+}$. This property ensures the uniqueness results for the equations (1.1) and (1.2), relying on the previous contributions [BSOT] and [DS].
    This unified approach allow us to compare, in the last part of this paper, the solutions of (1.1) and (1.2) by means of a doubling variables procedure.
  • (2)
    Hopf Lax Formula.
    With the aim of providing a self-contained existence result, that justifies our notion of viscosity solution, we consider, see Subsection 3.3, the function
    $U(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{d})}\left\{\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\right\},$
    where $\mathcal{L}:\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$ is the relaxed Lagrangian defined by
    $\mathcal{L}(\gamma)=\int_{\mathbb{R}^{2d}}L(v)d\gamma(x,v),$
    and $L:\mathbb{R}^{d}\to\mathbb{R}$ is the Lagrangian associated to the Hamiltonian $H$. Here $\exp_{\mu}:\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to\mathbb{R}$ denotes the exponential map in the Wasserstein space.
    By showing that the above formula provides a representation of the viscosity solution of (1.1) with convex Hamiltonian $H(x,p,\mu)=H(p)$ and a not necessarily differentiable Lagrangian, we generalize the existence result established in [BSOT].
    A related Hopf-Lax representation formula was previously introduced in [GANGBO2019119, Section  5], where first-order time dependent HJEs are considered in the case the Hamiltonian $H$ is convex and satisfies a structural decreasing property, see equation (5.7) there. This condotion ensures a compatibility between their lifted Hamiltonian and the Wasserstein geometry. In that setting, the Hopf-Lax representation, see for instance [GANGBO2019119, Section 5.4], is established in the quadratic case, relying crucially on the aforementioned decreasing property satisfied by the Hamiltonian. By contrast, our result does not require any such structural decreasing assumption of $H$.
  • (3)
    Geodesic Convexity and Idiosyncratic Operator. In Section 4, we inspection the action of the idiosyncratic operator on geodesic convex functions. To this end, we prove a connection between geodesic convexity and the convexity of the flat derivative w.r.t. the $\mathbb{R}^{d}$ variable. Indeed, we prove the following convexity result on mixtures of measures, which, to the best of our knowledge, is new in the literature.
    Lemma 4.6. Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continuous $\Lambda$-geodesically convex function. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then $\forall h\in[0,1]$ the function νF((1-h)μ+ hν) is $h\Lambda$-geodesically convex.
    The main difficulty in proving this result is that, even if $t\mapsto\nu_{t}$ is a $W_{2}$-geodesic, the curve $(1-h)\mu+h\nu_{t}$ is not. Nevertheless, this curve is still induced by transport coupling, so that using [CavSavSod23, Theorem 9.1], where the equivalence between geodesic convexity and total convexity is shown, we obtain a convex like inequality along this curve, which yields the desired conclusion. We remark that the dimensional restriction $d\geq 2$ comes from [CavSavSod23, Theorem 9.1]. Relying on this result, we do not know if our result can be extended to $d=1$.
    As a byproduct of the previous lemma, we obtain the following result
    Corollary 4.10. Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continous $\Lambda$-geodesically convex function. If $\mathcal{F}$ is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ then $x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is convex in $\mathrm{supp}(\mu)$.
    Here we denoted by $\mathcal{D}_{\mu}$ the flat derivative operator at $\mu$. The proof relies on the well known representation formula
    $\mathcal{D}_{\mu}\mathcal{F}(\mu,x):=\lim_{h\to 0}\frac{\mathcal{F}((1-h)\mu+h\delta_{x})-\mathcal{F}(\mu)}{h}\quad(\mu,x)\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathrm{supp}(\mu).$
    The above results are crucial to prove an estimate for the weak action of the idiosyncratic operator introduced in [DS].
    Proposition 4.5. Suppose $d\geq 2$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu^{-1}\in L^{\infty}_{\text{loc}}$ and $\mu\in W^{1,\infty}_{\text{loc}}$. Let $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}$ be a $\Lambda$-geodesically concave functional that is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
    $\int_{\mathbb{R}^{d}}\big<\nabla\log\mu,\nabla\mathcal{D}_{\mu}\mathcal{F}(\mu,x)\big>d\mu\leq d\Lambda.$
  • (4)
    Vanishing viscosity limit.
    The final result constitutes the main motivation of this article. Namely, showing the convergence of $U^{\varepsilon}$, the viscosity solution of (1.2), with $a(\mu,x)=\text{Id}_{d\times d}$, towards $U$ viscosity solution of (1.1), providing an optimal rate of convergence.
    Theorem 5.1. Under Assumption 2.15, let $U^{\varepsilon}$ and $U$ be bounded viscosity solutions of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}$, and (1.1), respectively. Assume that $U^{\varepsilon}(T,\mu)=U(T,\mu)=\mathcal{G}(\mu)$. Moreover, suppose $U^{\varepsilon}$ to be $W_{1}$-Lipschitz continuous, while $U$ to be $W_{2}$-Lipschitz continuous, Lipschitz in time, and continuous w.r.t. the weak convergence. Then there exists a constant $C\geq 0$ depending only on the data such that
    $\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}|U^{\varepsilon}(t,\mu)-U(t,\mu)|\leq C\sqrt{\varepsilon}.$
    We begin by briefly discussing the assumptions. The boundedness requirement on the solutions can be relaxed by introducing a suitable growth control; however, we decided not to focus on this technical issue. The main restriction of our result is the further continuity assumption w.r.t. the weak convergence of measures that we briefly discuss here: the space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is not locally compact, see [AGS, Remark  7.1.9], and this usually constitutes an issue in any optimization procedure, as in the doubling variables argument. However, $\mathcal{P}_{2}(\mathbb{R}^{d})$ is locally compact w.r.t. the weak convergence. We therefore impose further regularity w.r.t. this convergence in order to properly optimize. See the discussion in [bertucci2025doublingvariablestechniqueorder, Section  4.6] where tentative approaches to solve this important problem are also proposed.
    We now heuristically outline the main arguments in the lighter time-independent case, since the core argument is completely analogous to that of the evolutive case. Note that, to derive the vanishing-viscosity limit, we do not rely on possible explicit representations of the solution, as is often done in the control-theoretic setting. Instead, we adopt a purely PDE-based approach, following methods already established in the finite-dimensional case, see the monograph [Bardi1997OptimalCA, Section VI]. The main motivation for this choice is that this approach may also be useful in more general settings, such as the broader metric framework considered in [CKTComp].
    As in the finite dimensional case, we perform a doubling variables argument penalizing the difference between the two solutions:
    $(\mu,\nu)\mapsto U^{\varepsilon}(\mu)-U(\nu)-\frac{1}{2\alpha}W_{2}^{2}(\mu,\nu)-\delta\mathcal{E}(\mu)+\mbox{localizing term},$
    where $\mathcal{E}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is the entropy functional relative to the $d$-dimensional Lesbesgue measure $\lambda_{d}$. Thanks to this penalization, we obtain a maximum $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ (here we use the compactness given by the continuity w.r.t. the weak convergence of the two solutions) at which we can exploit the joint superdifferentiability of $W^{2}_{2}(\cdot,\cdot)$ at $(\overline{\mu},\overline{\nu})$. As usual, this procedure gives elements in the subdifferential of $U^{\varepsilon}-\delta\mathcal{E}$ at $\overline{\mu}$ and in the superdifferential of $U$ at $\overline{\nu}$, which can be used to apply the viscosity solution property to $U^{\varepsilon}$ and $U$. Let us also stress that the maximization gives $\mathcal{I}(\overline{\mu})<\infty$. The viscosity solution property yields
    $U^{\varepsilon}(\overline{\mu})-U(\overline{\nu})\leq C\alpha+K(\varepsilon)\delta-\underbrace{\varepsilon\int_{\mathbb{R}^{d}}\big<\nabla\log\overline{\mu}(x),\frac{x-T_{\overline{\mu}}^{\overline{\nu}}(x)}{\alpha}\big>d\overline{\mu}}_{I_{\varepsilon}}(x),$
    where the constant $C$ depends only on the local Lipschitz regularity of $H$. The term $C\alpha$ is obtained similarly in the proof of the comparison principle presented in [DS]. The main issue in our case, is the dependence of the constant $K(\varepsilon)$ on the diffusion, which quantitatively compensates the fact that we considered an element of the superdifferential of $U^{\varepsilon}-\delta\mathcal{E}$, and not merely of $U^{\varepsilon}$. Indeed, this constant is expected in general to blowup, see the formal computation in [DS, Section  2.3]. Therefore, we need a bound on $I_{\varepsilon}$ that does not depend on $\delta$. Indeed, applying a naive Cauchy Scwhartz inequality, one can prove that
    $I_{\varepsilon}\leq\mathrm{Lip}(U^{\varepsilon})\Big(\frac{\varepsilon}{\delta}+\varepsilon\alpha\Big)\leq C\Big(\frac{\varepsilon}{\delta}+\varepsilon\alpha\Big).$
    Optimizing w.r.t $\delta$ yelds
    $U^{\varepsilon}(\overline{\mu})-U(\overline{\nu})\leq C\alpha+\sqrt{K(\varepsilon)\varepsilon},$
    whose RHS does not, a priori, vanish as $\varepsilon\to 0.$ We therefore need a sharper bound on $I_{\varepsilon}$ that does not depend on $\delta$, namely
    $I_{\varepsilon}=-\varepsilon\int_{\mathbb{R}^{d}}\big<\nabla\log\overline{\mu}(x),\frac{x-T^{\overline{\nu}}_{\overline{\mu}}(x)}{\alpha}\big>d\overline{\mu}(x)\leq d.$
    as proven in Lemma 4.3, where the regularity assumptions therein are satisfied by the optimizer $\overline{\mu}\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$. This estimate can be regarded as the Wasserstein analogue of the finite-dimensional identity $\varepsilon\Delta_{x}\frac{|y-x|^{2}}{2\alpha}=\frac{d\varepsilon}{\alpha}$ which plays a crucial role in the original proof in [Bardi1997OptimalCA].
    The conclusion then follows as in the finite dimensional argument and, being the proof symmetric, we also obtain the converse inequality, completing the convergence result.
    We then conclude the analysis showing that the one side rate of convergence (from above) can be improved from $\sqrt{\varepsilon}$ to $\varepsilon$, provided $U^{\varepsilon}$ is a classical solution such that $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}\leq C$ uniformly (in the support of $\mu$). In particular, thanks to the analysis on geodesically concave functions, carried out in Section 4, this rate is achieved by $C^{2}$ geodesically $\lambda$-concave functions, at least for $d\geq 2$.
    Finally, the above rates are optimal since they are optimal111w.r.t. the assumptions on the Hamiltonian. See [cirant2025convergenceratesvanishingviscosity, chaintron2025optimalrateconvergencevanishing] for the optimal rate in the convex regime. in the finite dimensional case that embeds in this mean field framework.
1.3. Structure of the paper.
In Section 2, we fix the notation, we provide the main assumptions we are going to work with, and recall some important results from the literature that will be useful for our work. Moreover, we investigate some links between various notions of differentials. As a consequence, we will determine one meaningful differential to treat in the viscosity sense both the equations (1.2) and (1.2).
In Section 3, we prove in an intrinsic way a Hopf-Lax representation of the solution of (1.1). In Section 4, we derive estimates and lemmata that will be useful in Section 5, where we finally establish the vanishing viscosity result.
2. The Wasserstein Space and the Related Subdifferential Calculus
2.1. The Wasserstein space
Our ambient space will be the space of probability measures on the $d$-dimensional Euclidean space $\mathbb{R}^{d}$ denoted as $\mathcal{P}(\mathbb{R}^{d})$. The measure $\lambda_{d}$ will stand for the Lesbesgue measure in $\mathbb{R}^{d}$, and in the integration we will often use $dx$ rather than $\lambda_{d}$.
Given $\mu\in\mathcal{P}(\mathbb{R}^{d})$, we will denote by
$\mathcal{M}_{2}(\mu)$ $:=\int_{\mathbb{R}^{d}}|x|^{2}d\mu,$
$\mathcal{P}_{2}(\mathbb{R}^{d})$ $:=\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\mathcal{M}_{2}(\mu)<\infty\},$
$\mathcal{P}_{2,ac}(\mathbb{R}^{d})$ $:=\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}):\mu<<\lambda_{d}\}$
the second moment of the measure $\mu$ and the $L^{2}$-Wasserstein space, and the space of absolutely continuous measures with respect to $\lambda_{d}$, respectively.
The Wasserstein space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a complete and separable geodesic metric space when equipped with the Wasserstein distance $W_{2}:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,\infty)$, defined for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ as
(2.1) $W_{2}(\mu,\nu):=\min_{\gamma\in\Gamma(\mu,\nu)}W^{\gamma}_{2}(\mu,\nu),$
where
$\Gamma(\mu,\nu):=\left\{\gamma\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d}):{(\pi_{1})}_{\#}\gamma=\mu,\,{(\pi_{2})}_{\#}\gamma=\nu\right\}$
is the set of couplings with marginals $\mu$ and $\nu$, $\pi_{i}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the standard projection operator onto the $i$-th coordinate, $i\in\{1,2\}$, and, for a coupling $\gamma\in\Gamma(\mu,\nu),$
(2.2) $W^{\gamma}_{2}(\mu,\nu):=\left(\int_{\mathbb{R}^{2d}}|x-y|^{2}d\gamma(x,y)\right)^{\frac{1}{2}}$
is the transport cost from $\mu$ to $\nu$ associated to $\gamma$. For further details, see [OTVillani, Chapter 4, Chapter 6].
We denote by $\Gamma_{0}(\mu,\nu)$ the set of optimal couplings between $\mu$ and $\nu$ for the distance $W_{2}$, i.e., any coupling that realizes the minimum cost in (2.1). In general, we will say that a plan $\gamma\in\mathcal{P}(\mathbb{R}^{2d})$ is optimal if $\gamma\in\Gamma_{0}((\pi_{1})_{\#}\gamma,(\pi_{2})_{\#}\gamma)$, i.e. $\gamma$ is optimal w.r.t. its marginals.
Moreover, for ${\gamma}\in\mathcal{P}(\mathbb{R}^{2d}),\nu\in\mathcal{P}(\mathbb{R}^{d})$, the sets
$\Gamma({\gamma},\nu):=\{\theta\in\mathcal{P}(\mathbb{R}^{3d}):(\pi_{1},\pi_{2})_{\#}\theta=\gamma,\ (\pi_{1},\pi_{3})_{\#}\theta\in\Gamma({(\pi_{1})}_{\#}\gamma,\nu)\}$
$\Gamma_{0}({\gamma},\nu):=\{\theta\in\mathcal{P}(\mathbb{R}^{3d}):(\pi_{1},\pi_{2})_{\#}\theta=\gamma,\ (\pi_{1},\pi_{3})_{\#}\theta\in\Gamma_{0}({(\pi_{1})}_{\#}\gamma,\nu)\}$
will denote the set of (resp. optimal) couplings between the plan $\gamma$ and the measure $\nu$. Here $\pi_{i}:\mathbb{R}^{3d}\ \to\mathbb{R}^{d}$ is the standard projection operator onto the $i$-th coordinate, $i\in\{1,2,3\}$, coherently with the notation used for the projection from $\mathbb{R}^{2d}$ to $\mathbb{R}^{d}$.
To any $\gamma\in\mathcal{P}(\mathbb{R}^{2d})$ we associate its disintegration w.r.t. the first marginal $\mu:=(\pi_{1})_{\#}\gamma$: i.e. the $\mu$-a.e. uniquely determined measure $\gamma_{x}(dy)\in\mathcal{P}(\mathbb{R}^{d})$ s.t. $\gamma=\gamma_{x}(dy)\mu(dx)$. The disintegration easily generalizes to the second marginal and to every marginal of measures $\theta\in\mathcal{P}(\mathbb{R}^{3d})$. See [AGS, Theorem 5.3.1].
Let $\mathcal{F}\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a function. We call
$\mathrm{Lip}(\mathcal{F};W_{2}):=\inf_{\mu\not=\nu}\frac{|\mathcal{F}(\mu)-\mathcal{F}(\nu)|}{W_{2}(\mu,\nu)}$
the Lipschitz constant of $\mathcal{F}$ w.r.t. the $W_{2}$ distance.
For future use, we mention that for any $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, as a simple consequence of the Cauchy-Schwartz inequality we have the following inequalities
$|\mathcal{M}_{2}(\mu)-\mathcal{M}_{2}(\nu)|\leq W_{2}(\mu,\nu)\left(\mathcal{M}_{2}(\mu)^{\frac{1}{2}}+\mathcal{M}_{2}(\nu)^{\frac{1}{2}}\right).$
In particular, we have the following locally Lipschitz bound for second moments
(2.3) $|\mathcal{M}_{2}(\mu)-\mathcal{M}_{2}(\nu)|\leq\frac{1}{2}W_{2}(\mu,\nu)\left(2+\mathcal{M}_{2}(\mu)+\mathcal{M}_{2}(\nu)\right).$
Indeed, given $\sigma\in\Gamma_{0}(\mu,\nu)$, by standard algebra, we have
$\mathcal{M}_{2}(\mu)-\mathcal{M}_{2}(\nu)$ $=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(x-y)\cdot(x+y)d\sigma(x,y)$
$\leq\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{2}d\sigma(x,y)\left){}^{\frac{1}{2}}\right(\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x+y|^{2}d\sigma(x,y)\right)^{\frac{1}{2}}$
$\leq W_{2}(\mu,\nu)\left((\int_{\mathbb{R}^{d}}|x|^{2}d\mu)^{\frac{1}{2}}+(\int_{\mathbb{R}^{d}}|y|^{2}d\nu(y))^{\frac{1}{2}}\right).$
2.2. Tangent spaces at a measure
We survey here the definitions of various notions of tangent space at a measure, as described in [AGS, Appendix], the definition of tangent bundle and their links. Thanks to this description, we can unify the notions of viscosity solution in the deterministic case and in the presence of idiosyncratic noise.
Definition 2.1.
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. The (Monge) tangent space at $\mu$ is defined as the set
$\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}):=\overline{\left\{\varepsilon(r-\mbox{Id})\colon r:\mathbb{R}^{d}\to\mathbb{R}^{d}\mbox{ Borel measurable},(\mbox{Id}\times r)_{\#}\mu\text{ is optimal, for }\varepsilon>0\right\}}^{L^{2}_{\mu}(\mathbb{R}^{d})},$
where $L^{2}_{\mu}(\mathbb{R}^{d})$ is the $L^{2}$ space of $\mu$-measurable $\mathbb{R}^{d}$-valued maps.
Given $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, let us call
$\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}):=\left\{\gamma\in\mathcal{P}_{2}{(\mathbb{R}^{2d})},(\pi_{1})_{\#}\gamma=\mu\right\},$
the set of plans in $\mathcal{P}_{2}{(\mathbb{R}^{2d})}$ having $\mu$ as first marginal, and for $\gamma,\xi\in\mathcal{P}_{2,\mu}({\mathbb{R}^{2d}})$
$\Gamma_{\mu}(\gamma,\xi):=\{\theta\in\mathcal{P}(\mathbb{R}^{3d}),(\pi_{1,2})_{\#}\theta=\gamma,(\pi_{1,3})_{\#}\theta=\xi\},$
the set of plans in $\mathcal{P}{(\mathbb{R}^{3d})}$, couplings in $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$. Here, $\pi_{1,j}:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}\times\mathbb{R}^{d}$ is the projection on the first and the $j$-th coordinate $\pi_{1,j}:=(\pi_{1},\pi_{j})$, $j=2,3$. To shorten the notation we will write $\gamma^{\mu}$ instead of $\gamma$, when the latter belongs to $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$.
The set $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ is metrized by the distance $W_{\mu}\colon\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\times\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to[0,\infty)$ defined in [AGS, Equation  (12.4.10)], as
(2.4) $W^{2}_{\mu}(\gamma^{\mu}_{1},\gamma^{\mu}_{2}):=\min_{\theta\in\Gamma_{\mu}(\gamma_{1}^{\mu},\gamma_{2}^{\mu})}\left\{\int_{\mathbb{R}^{3d}}|y-z|^{2}d\theta(x,y,z)\right\}.$
Note that the convergence induced by $W_{\mu}$ on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ coincides with that induced by the natural extension of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R}^{2d})$ restricted to $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$. [^2] To see this: Let us consider the natural extension of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R}^{2d})$ defined as $W^{2}_{2}(\gamma_{1},\gamma_{2})=\min_{\theta\in\Gamma(\gamma_{1},\gamma_{2})}\int_{\mathbb{R}^{4d}}(|x_{2}-x_{1}|^{2}+|z_{1}-z_{2}|^{2})d\theta(x_{1},z_{1};x_{2},z_{2})$ for $\gamma_{1},\gamma_{2}\in\mathcal{P}_{2}(\mathbb{R}^{2d})$. Here $\Gamma(\gamma_{1},\gamma_{2})=\big\{\sigma\in\mathcal{P}(\mathbb{R}^{2d}\times\mathbb{R}^{2d}):(\pi_{i})_{\#}\sigma=\gamma_{i},i={1,2}\big\},$ where $\pi_{i}:\mathbb{R}^{2d}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d},i=1,2$ are the natural projections to the first and second factor, respectively. We note that $\int_{\mathbb{R}^{4d}}(|x_{2}-x_{1}|^{2}+|z_{1}-z_{2}|^{2})d\theta(x_{1},z_{1};x_{2},z_{2})\geq\int_{\mathbb{R}^{4d}}|z_{1}-z_{2}|^{2}d\theta(x_{1},z_{1};x_{2},z_{2})$ $\forall\theta\in\Gamma(\gamma_{1},\gamma_{2})\subset\mathcal{P}(\mathbb{R}^{4d})$, and the equality holds iff $x_{1}=x_{2},\,\theta$-a.e., i.e. $\gamma_{1},\gamma_{2}$ shares the same first marginal and $(\pi_{1},\pi_{3})_{\#}\theta=(\text{Id}\times\mathrm{Id})_{\#}\mu$. In particular, $\theta$ can be identified as a plan in $\mathcal{P}(\mathbb{R}^{3d})$ belonging to $\Gamma_{\mu}(\gamma_{1},\gamma_{2})$, and $W_{\mu}$ coincides with the restriction of the $W_{2}$ to $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$.
Definition 2.2.
The (Kantorovich) geometric tangent space at $\mu$ is defined as:
$\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}):=\overline{\left\{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}):(\pi_{1},\pi_{1}+\varepsilon\pi_{2})_{\#}\gamma\text{ is optimal, for some }\varepsilon>0\right\}}^{W_{\mu}(\cdot,\cdot)}$
In $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$, it is naturally defined an exponential map
$\exp_{\mu}:$ $\ \mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to\mathcal{P}_{2}(\mathbb{R}^{d})$
$\ \gamma\longmapsto(\pi_{1}+\pi_{2})_{\#}\gamma.$
We remark that, as in [GigliPhd, Definition  4.18], we prefer to define the exponential on the whole $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ rather than on its closed subset $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ for reasons that will become clear later in our discussion about differentials. Moreover, let us note that $\exp_{\mu}$ is surjective, i.e., more precisely, for all $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists an optimal velocity $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\exp_{\mu}(\gamma)=\nu$.
Following [GigliPhd, Definition  4.13], given $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we can also introduce a notion of “scalar product” on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ as
(2.5) $\big<\gamma_{1}^{\mu},\gamma_{2}^{\mu}\big>_{\mu}:=\max_{\theta\in\Gamma_{\mu}(\gamma_{1}^{\mu},\gamma_{2}^{\mu})}\big<\gamma_{1}^{\mu},\gamma_{2}^{\mu}\big>_{\theta,\mu},$
where for $\theta\in\Gamma_{\mu}(\gamma_{1}^{\mu},\gamma_{2}^{\mu})$, $\big<\gamma_{1}^{\mu},\gamma_{2}^{\mu}\big>_{\theta,\mu}:=\int_{\mathbb{R}^{3d}}\big<y,z\big>d\theta(x,y,z)$.
Note that the extrema of (2.5) and (2.4) coincide because of the identity $|y-z|^{2}=|y|^{2}+|z|^{2}-2\big<y,z\big>,\quad\forall y,z\in\mathbb{R}^{d}$. Moreover, whenever either $\gamma_{1}^{\mu}$ or $\gamma_{2}^{\mu}$ is induced by a map, say $\gamma_{2}^{\mu}=(\mathrm{Id}\times p)_{\#}\mu$ with $p\in L^{2}_{\mu}(\mathbb{R}^{d})$, being $\Gamma_{\mu}(\gamma_{1}^{\mu},\gamma_{2}^{\mu})$ a singleton, we have
$\big<\gamma_{1}^{\mu},\gamma_{2}^{\mu}\big>_{\mu}=\int\big<y,p(x)\big>d\gamma_{1}^{\mu}(x,y).$
As a consequence one can also introduce a Wasserstein norm in $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$, defined for $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ as
(2.6) $\|\xi\|_{\mu}:=\sqrt{\big<\xi,\xi\big>_{\mu}}=W_{\mu}(\xi,(\mathrm{Id}\times 0)_{\#}\mu).$
In particular, the norm is continuous w.r.t. the convergence induced by $W_{\mu}$. Moreover, observe that
$W_{2}(\mu,\exp_{\mu}(\xi))\leq\|\xi\|_{\mu},\quad\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
where the equality holds if $\xi\in\Gamma_{0}(\mu,\exp_{\mu}(\xi)).$
By [GigliPhd, Proposition  4.21], we have that, for $\gamma_{1},\gamma_{2},\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$,
$|\big<\gamma_{1},\xi\big>_{\mu}-\big<\gamma_{2},\xi\big>_{\mu}|\leq W_{\mu}(\gamma_{1},\gamma_{2})\|\xi\|_{\mu}.$
In other words, the scalar product is Lipschitz continuous w.r.t. $W_{\mu}$. We also remark, see [GigliPhd, Remark  4.28], that in general $\max_{\theta\in\Gamma_{\mu,\nu}(\gamma,\xi)}\big<\gamma,\xi\big>_{\theta,\mu}>\min_{\theta\in\Gamma_{\mu,\nu}(\gamma,\xi)}\big<\gamma,\xi\big>_{\theta,\mu}.$ However, the equality holds if either $\gamma$ or $\xi$ is induced by a map. Lastly, it is defined a multiplication operation, see in [GigliPhd, Proposition  4.22 &\& Proposition  4.24]: given $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$\mathbb{R}\ni\lambda\mapsto\lambda\cdot\gamma:=(\pi_{1},\lambda\pi_{2})_{\#}\gamma.$
If $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, it is immediate to see that $\lambda\cdot\gamma\in\mathrm{Tan}_{\mu}^{K}{\mathcal{P}}_{2}(\mathbb{R}^{d})$, for every $\lambda\geq 0$. This invariance is still true for $\lambda<0$, see [GigliPhd, Proposition  4.29]. To lighten the notation we will often write $\gamma$ for $1\cdot\gamma$. We also remark that the product $\big<\cdot,\cdot\big>_{\mu}$ enjoys a one side Cauchy-Schwarz inequality: Indeed, for every $\gamma^{1},\gamma^{2}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$\big<\gamma^{1},\gamma^{2}\big>_{\mu}$ $=\max_{\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})}\int\big<w_{1},w_{2}\big>d\theta(x,w_{1},w_{2})$
$\leq\max_{\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})}\Big(\int|w_{1}|^{2}d\theta(x,w_{1},w_{2})\Big)^{\frac{1}{2}}\Big(\int|w_{2}|^{2}d\theta(x,w_{1},w_{2})\Big)^{\frac{1}{2}}$
$=\Big(\int|w_{1}|^{2}d\gamma^{1}(x,w_{1})\Big)^{\frac{1}{2}}\Big(\int|w_{2}|^{2}d\gamma^{2}(x,w_{2})\Big)^{\frac{1}{2}}$
$=\|\gamma^{1}\|_{\mu}\|\gamma^{2}\|_{\mu}.$
Moreover, the equality holds iff there exists $\theta\in\Gamma_{\mu}(\gamma^{1},\gamma^{2})$ optimizer for $\big<\cdot,\cdot\big>_{\mu}$ s.t. $\big<w_{1},w_{2}\big>=|w_{1}||w_{2}|$ holds $\theta$-a.e., i.e. $\gamma^{1}=\lambda\gamma^{2}$, with $\lambda\geq 0$.
As in [GigliPhd, Proposition  4.30], we define a projection operator on $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{2d})$:
$\text{P}:\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),$
that at each element $\gamma$ associates the unique plan $\text{P}(\gamma)\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ that minimizes the distance between $\gamma$ and the geometric tangent space $\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. In particular, see [GigliPhd, Corollary  4.34],
(2.7) $\big<\text{P}(\gamma),\eta\big>_{\mu}=\big<\gamma,\eta\big>_{\mu}\quad\forall\eta\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\quad\text{and}\quad\|\text{P}(\gamma)\|_{\mu}\leq\|\gamma\|_{\mu},\quad\forall\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}).$
In order to understand the link between the two different notions of tangent space, let us consider the barycentric projection (see [AGS, Definition  5.4.2.]). For a fixed $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the barycentric projection map is defined as
$\text{Pr}_{\mu}\colon\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ $\to L^{2}_{\mu}(\mathbb{R}^{d}),$
$\gamma$ $\mapsto\overline{\gamma},$
where, given $\gamma(dx,dy)=d\gamma_{x}(y)d\mu(x)$, the disintegration of $\gamma$ with respect to its first marginal $\mu=(\pi_{1})_{\#}\gamma$, the image $\overline{\gamma}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is defined as :
$\overline{\gamma}(x)=\int_{\mathbb{R}^{d}}y\,d\gamma_{x}(y).$
Note that Jensen inequality ensures the wellposedness of the above map, indeed $\overline{\gamma}\in L^{2}_{\mu}(\mathbb{R}^{d})$. To be more general, we can define the projection map [^3] In a probabilistic framework, the barycentric projection of the law of a random vector $(X,Y)$ corresponds to the law of the conditional expectation $\mathbb{E}[Y|X]$. on the whole $\mathcal{P}_{2}(\mathbb{R}^{2d})$, as
$\text{Pr}(\gamma):=\text{Pr}_{(\pi_{1})_{\#}\gamma}(\gamma),\quad\forall\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}).$
It turns out that, for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the Monge tangent space can be seen as the image of the restriction of the barycentric projection to the geometric tangent space. Moreover, it is not difficult to show that Pr is onto and a 11-contraction (see [AGS, Definition 12.4.3], and the subsequent discussion). Namely
(2.8) $\text{Pr}(\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}))$ $=\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),$
(2.9) $\|\text{Pr}(\gamma_{1})-\text{Pr}(\gamma_{2})\|_{L^{2}_{\mu}(\mathbb{R}^{d})}$ $\leq W_{\mu}(\gamma_{1},\gamma_{2})\quad\forall\gamma_{1},\gamma_{2}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}).$
When $\mu$ is absolutely continuous, we can be more precise: this projection is a one-to-one isometry w.r.t. the distances previously introduced, as shown in [AGS, Theorem  12.4.4]. Thus
(2.10) $(\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),W_{\mu})\simeq_{\text{Pr}}(\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),\|\cdot\|_{L^{2}_{\mu}}).$
Let us note that, using the characterization of optimal maps for absolutely continuous measures, we also have the following alternative description for the tangent space (see [AGS, Definition  8.4.1 &\& Theorem  8.5.1]):
(2.11) $\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})=\overline{\left\{\nabla\phi\colon\phi\in C^{\infty}_{c}(\mathbb{R}^{d})\right\}}^{L^{2}_{\mu}(\mathbb{R}^{d})}.$
In particular, at an absolutely continuous measure $\mu$ the scalar product (2.5) is the standard scalar product of $L_{\mu}^{2}(\mathbb{R}^{d})$ restricted to $\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$.
2.3. Subdifferential calculus in the Wasserstein space.
2.3.1. First variation and Flat structure
We now briefly recap the principal notions of differential calculus in Wasserstein spaces. For more details, see [AGS, Section  10.3] and [CarmonaDelarue, Chapter 5]. At the end of the discussion, we will select a unique differential that can be used for equations (1.1) and (1.2).
We first observe that the Wasserstein space can be seen as a convex subset of the vector space of finite Borel measures $\mathcal{M}(\mathbb{R}^{d})$. We prefer to give a more general definition inspired by the definitions of first derivative over $\mathcal{P}_{2}(\mathbb{R}^{d})$ contained in [PierreCaNotes] and [santambrogio2015optimal]. The growth assumption is taken by [CarmonaDelarue, Chapter  5].
Definition 2.3.
Let $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$. We say that $U$ is flat differentiable at $\mu$, if there exists a Borel measurable map with quadratic growth denoted
$\mathcal{D}_{\mu}U(\mu,\cdot):\mathbb{R}^{d}\to\mathbb{R},$
s.t., $\forall\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, it holds
$\lim_{h\to 0^{+}}\frac{U((1-h)\mu+h\nu)-U(\mu)}{h}=\int_{\mathbb{R}^{d}}\mathcal{D}_{\mu}U(\mu,x)(\nu-\mu)(dx)\,.$
We say that $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is $C^{1}$ (resp. partially $C^{2}$) if $\mathcal{D}_{\mu}U:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}$ is jointly continuous with uniform quadratic growth (there exists $\frac{\partial}{\partial x}\mathcal{D}_{\mu}U$ that is jointly continuous and with uniform quadratic growth).
Since the convex interpolation is restricted to be in $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\int\text{cost}(\nu-\mu)dy=0$, in supp($\mu$), the flat derivative $\mathcal{D}_{\mu}U$ is defined up to a constant in depending only on $\mu$. We select this constant s.t. $\int_{\mathbb{R}^{d}}\mathcal{D}_{\mu}U(\mu,x)d\mu=0.$ Thanks to this choice of the constant, we have the following representation formula: if $U$ is flat differentiable at $\mu$, then
(2.12) $\mathcal{D}_{\mu}U(\mu,y)=\lim_{h\to 0^{+}}\frac{1}{h}\left(U\left((1-h)\mu+h\delta_{y}\right)-U(\mu)\right),\quad\forall y\in\text{supp}(\mu).$
2.3.2. Wasserstein subdifferential calculus
Various notions of subdifferentiability have been introduced in the literature. Let us start with two classical notions that are given with two different points of view. The first one sees the elements of the differential as plans while the second one as vector valued functions. The two notions can be related via the barycentric projection map introducing an intermediate notion of differential.
In the following, we will provide only the definition of subdifferential of a proper and lower semicontinuous function since the superdifferential of a proper and upper semicontinuous function $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to[-\infty,\infty)$ can be obtained setting $\partial^{+}_{\cdot}\phi(\mu):=\big\{-1\cdot\gamma:\gamma\in\partial_{\cdot}^{-}(-\phi)(\mu)\big\}$ in the definitions 2.4 2.6, 2.7 and 2.9. As already remarked, this operation leaves invariant the geometric tangent space. Whenever $\partial^{-}_{\cdot}(-\phi)(\mu)\subset L^{2}_{\mu}(\mathbb{R}^{d})$, as in Definitions 2.5 and 2.8, we simply consider $\partial^{+}_{\cdot}\phi(\mu)=-\partial^{-}_{\cdot}(-\phi)(\mu)$. We denote by $\text{Dom}(\phi):=\Big\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}):\phi(\mu)<\infty\Big\}$
In what follows, we introduce the various notions of differentials within the formalism established above. For each definition, we explicitly indicate the corresponding reference.
Definition 2.4 (See [GigliPhd], Definition 5.14).
Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the Kantorovich subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{K}\phi(\mu)$, if for all $\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$
$\phi(\exp_{\mu}(\xi))-\phi(\mu)\geq-\big<-\gamma,\xi\big>_{\mu}+o(\|\xi\|_{\mu}).$
Note that in [GigliPhd, Definition 5.14], the author additionally requires that $\partial^{-}_{K}\phi(\mu)\subset\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. In the present work, we choose not to impose this restriction at the level of the definition, in order to allow for a more refined comparison with other notions.
Definition 2.5.
See [GANGBO2019119], Definition 3.1 ] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $p\in L^{2}_{\mu}(\mathbb{R}^{d})$ belongs to the Monge subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{M}\phi(\mu)$, if for all $\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$
$\phi(\exp_{\mu}(\xi))-\phi(\mu)\geq\int_{\mathbb{R}^{2d}}\langle p(x),z\rangle\,d\xi(x,z)+o(\|\xi\|_{\mu}).$
In order to formalize the link between the previous two classical notions, let us introduce the following intermediate definition, given in the spirit of Defintion 2.9 below, introduced by Bertucci in [BSOT].
Definition 2.6 (Intermediate Definition).
Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the intermediate Kantorovich subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{I}\phi(\mu)$, if for all $\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$
$\phi(\exp_{\mu}(\xi))-\phi(\mu)\geq\int_{\mathbb{R}^{2d}}\big<y,z\big>\gamma_{x}(dy)d\xi(x,z)+o(\|\xi\|_{\mu}),$
where $\gamma_{x}$ is the disintegration of $\gamma$ with respect to its first marginal $\mu$, i.e. $\gamma=\gamma_{x}(dy)\mu(dx)$.
Let us note that
$\partial^{-}_{I}\phi(\mu)\subseteq\partial^{-}_{K}\phi(\mu)\quad\text{and}\quad\text{Pr}(\partial^{-}_{I}\phi(\mu))=\partial^{-}_{M}\phi(\mu).$
Moreover, intersecting the subdifferential with the geometric tangent space, due to the properties of the barycentric projection map on the geometric tangent space, see equation (2.8), we have
$\text{Pr}(\partial^{-}_{I}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}))=\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}).$
Suppose $\mu$ is absolutely continuous, then $\mathrm{Tan}^{K}_{\mu}\mathcal{P}(\mathbb{R}^{d})$ can be isometrically identified with $\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\subset L^{2}_{\mu}(\mathbb{R}^{d})$. Therefore, the three inequalities above can be, equivalently, tested on maps and read: for all $\xi\in L^{2}_{\mu}(\mathbb{R}^{d})$
$\phi(\exp_{\mu}((\mathrm{Id}\times\xi)_{\#}\mu))-\phi(\mu)\geq\int_{\mathbb{R}^{d}}\big<\Pr(\gamma)(x),\xi(x)\big>d\mu(x)+o(\|\xi\|_{L^{2}_{\mu}}),$
since in this case $\big<(\text{Id}\times\xi)_{\#}{\mu},\gamma\big>_{\mu}=\big<\xi,\Pr(\gamma)\big>_{L^{2}_{\mu}(\mathbb{R}^{d})}.$ Moreover, being the barycentric projection a one-to-one isometry, we have
$\partial^{-}_{K}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\simeq_{\text{Pr}}\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),$
thus establishing a strong connection between the two differentials in the case of absolutely continuous measures.
Note that the set $\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the subdifferential set defined in [DS].
The unique element in $\partial^{+}_{K}\phi(\mu)\cap\partial^{-}_{K}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, if any, is called the Wasserstein differential of $\phi$, see [GigliPhd, Definition  5.14]. In this case, we say that $\phi$ is differentiable at $\mu$ and $\partial_{\mu}\phi$ is the Wasserstein differential at $\mu$.
A stronger notion of subdifferentials can be obtained by requiring, in Definitions 2.4,2.5 and 2.6, that the corresponding inequalities hold for a larger class of perturbations, e.g. without assuming the measure vector field $\xi$ to lie in the tangent space. This restriction will select more stable elements in the subdifferentials, and it has already been used in the gradient flow theory developed in[AGS] to treat the Euler equation of the minima of the Yosida approximation of a given functional. We also remark that this strongest notion is linked with the so-called Hilbertian lift. See [BSOT, Remark  2.5 &\& Proposition  2.6] and the introduction in [GANGBO2019119].
Definition 2.7.
[See [AGS, Definition  10.3.1] Strong Fréchet Subdifferential] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the strong subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ and $\theta\in\Gamma_{\mu}(\gamma,\xi)$,
$\phi(\exp_{\mu}(\xi))-\phi(\mu)\geq\big<\xi,\gamma\big>_{\theta,\mu}+o(\|\xi\|_{\mu}).$
Definition 2.8.
Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $p\in L^{2}_{\mu}(\mathbb{R}^{d})$ belongs to the (strong-Monge)subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S,M}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$\phi(\exp_{\mu}(\xi))-\phi(\mu)\geq\int_{\mathbb{R}^{2d}}\big<p(x),z\big>d\xi(x,z)+o(\|\xi\|_{\mu}).$
Definition 2.9.
[See [BSOT], Definition 2.2] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the map strong subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S,I}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$\phi(\exp_{\mu}(\xi))-\phi(\mu)\geq\int_{\mathbb{R}^{3d}}\big<y,z\big>\gamma_{x}(dy)d\xi(x,z)+o(\|\xi\|_{\mu}),$
where $\gamma_{x}$ is the disintegration of $\gamma$ with respect to its first marginal $\mu$, i.e. $\gamma=\gamma_{x}(dy)\mu(dx)$.
The above definition was introduced in [BSOT] to prove the well-posedness of viscosity solutions of Hamilton-Jacobi equation in $\mathcal{P}(\mathbb{T}^{d})$.
Remark 2.10.
Let us note that w.r.t. [BSOT, Definition  2.2], we preferred here to describe the elements of the subdifferential as plans instead of using the equivalent identification of $\gamma$ with $\gamma_{x}$ $\mu$-a.e. determined by $\gamma(dx,dy)=\gamma_{x}(dy)\mu(dx)$. We will do the same for the notion of viscosity solutions adopted in the first-order case.
As for the previous definitions, let us note that
$\partial^{-}_{S}\phi(\mu)\subseteq\partial^{-}_{S,I}\phi(\mu),\quad\quad\text{Pr}(\partial^{-}_{S,I}\phi(\mu))=\partial^{-}_{S,M}\phi(\mu)\quad\text{ and }\quad\text{Pr}(\partial^{-}_{S}\phi(\mu))\subseteq\partial_{S,M}^{-}\phi(\mu)$
Moreover, we have
$\text{Pr}(\partial^{-}_{S,I}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}))=\partial^{-}_{S,M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),$
while the isometries
$\partial^{-}_{S,I}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\simeq_{\text{Pr}}\partial^{-}_{S,M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$
holds for absolutely continuous $\mu$.
Remark 2.11.
Let us note that
$\partial^{-}_{S}\phi(\mu)\subseteq\partial^{-}_{S,I}\phi(\mu)\subseteq\partial^{-}_{I}\phi(\mu)\subseteq\partial^{-}_{K}\phi(\mu),$
and
$\partial^{-}_{S,M}\phi(\mu)\subseteq\partial^{-}_{M}\phi(\mu).$
Obviously, the same inclusions holds, with suitable modifications, for the superdifferentials.
Lemma 2.12.
(Correspondent of [GANGBO2019119, Theorem  3.14] in our setting) Let $\phi:\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be a proper lower semicontinuous functional, and $\mu\in\text{Dom}(\phi)$. Then,
$\partial_{S,M}^{-}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ $=\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}).$
Moreover, when $\mu$ is absolutely continuous w.r.t. the Lesbsegue measure
(2.13) $\partial^{-}_{S}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\simeq_{\mathrm{Pr}}\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}).$
Proof.
In view of the inclusion $\partial^{-}_{S,M}\phi(\mu)\subseteq\partial^{-}_{M}\phi(\mu)$, it is sufficient to prove only the $\supseteq$ part. The proof is an adaptation of [Erbar2010, Lemma  3.2].
Suppose the the inclusion does not hold: There exists $p\in\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),\,\delta>0$ and a sequence $(\xi_{n})_{n}\subset\mathcal{P}_{2,\mu}(\mathbb{R}^{d})$ with $\|\xi_{n}\|_{\mu}=\varepsilon_{n}\to 0$ s.t.
$\phi(\exp_{\mu}(\xi_{n}))-\phi(\mu)-\int_{\mathbb{R}^{2d}}\big<p(x),z\big>d\xi_{n}(x,z)\leq-\delta\varepsilon_{n}.$
Choose $(\widetilde{\xi}_{n})_{n}\subset\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\forall n\in\mathbb{N}$
$\exp_{\mu}(\xi_{n})=\exp_{\mu}(\widetilde{\xi}_{n})$
$\|\widetilde{\xi}_{n}\|_{\mu}=W_{2}(\mu,\exp_{\mu}(\widetilde{\xi}_{n}))=W_{2}(\mu,\exp_{\mu}(\xi_{n}))\leq\|\xi_{n}\|_{\mu}=\varepsilon_{n}.$
In other words, $\widetilde{\xi}_{n}$ is an optimal velocity plan connecting $\mu$ and $\exp_{\mu}(\xi_{n})$. Since $p\in\partial^{-}_{M}\phi(\mu)$, there exists $N>0$ s.t.
$\phi(\exp_{\mu}(\xi_{n}))-\phi(\mu)-\int_{\mathbb{R}^{2d}}\big<p(x),z\big>d\widetilde{\xi}_{n}(x,z)\geq-\frac{1}{2}\delta\varepsilon_{n},\quad\forall n>N.$
By combining the two inequality we deduce that
$\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\text{Pr}(\widetilde{\xi}_{n})}{\varepsilon_{n}}\big>_{\mu}-\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\text{Pr}(\xi_{n})}{\varepsilon_{n}}\big>_{\mu}$ $=\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\widetilde{\xi}_{n}}{\varepsilon_{n}}\big>_{\mu}-\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\xi_{n}}{\varepsilon_{n}}\big>_{\mu}$
$\leq-\frac{\delta}{2}\varepsilon_{n}\quad\forall n>N,$
where we used the scalar product defined in (2.5) and its property along graphs. We now observe that
[…]
are both bounded by construction. We can then extract a, not relabeled, weak limit in $L^{2}_{\mu}(\mathbb{R}^{d})$ of the two sequences $\left(\frac{\text{Pr}(\widetilde{\xi}_{n})}{\varepsilon_{n}}\right)_{n},\left(\frac{\text{Pr}(\xi_{n})}{\varepsilon_{n}}\right)_{n}$, call them $\widetilde{\xi},{\xi}\in L^{2}_{\mu}(\mathbb{R}^{d})$, respectively. In particular,
(2.14) $\big<(\mathrm{Id}\times p)_{\#}\mu,\widetilde{\xi}\big>_{\mu}-\big<(\mathrm{Id}\times p)_{\#}\mu,\xi\big>_{\mu}=$
$=\lim_{n\to\infty}\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\text{Pr}(\xi_{n})}{\varepsilon_{n}}\big>_{\mu}-\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\text{Pr}(\widetilde{\xi}_{n})}{\varepsilon_{n}}\big>_{\mu}$
$=\lim_{n\to\infty}\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\xi_{n}}{\varepsilon_{n}}\big>_{\mu}-\big<(\mathrm{Id}\times p)_{\#}\mu,\frac{\widetilde{\xi}_{n}}{\varepsilon_{n}}\big>_{\mu}\leq-\frac{\delta}{2}.$
On the other hand, due to the construction of the plans $\xi_{n},\widetilde{\xi}_{n}$, we have that, for any $\alpha_{n}\in\Gamma_{\mu}(\xi_{n},\widetilde{\xi}_{n})$, the equality
$\int_{\mathbb{R}^{3d}}\Big(\psi(z_{0}-x)-\psi(z_{1}-x)\Big)d\alpha_{n}(x,z_{0},z_{1})=\int_{\mathbb{R}^{3d}}\psi(y)d\exp_{\mu}(\xi_{n})(y)-\int_{\mathbb{R}^{3d}}\psi(y)d\exp_{\mu}(\widetilde{\xi}_{n})(y)=0,$
holds $\forall\psi\in C^{\infty}_{c}(\mathbb{R}^{d})$. Moreover, given $\psi\in C^{\infty}_{c}(\mathbb{R}^{d})$, by a Taylor expansion of the function and its global semiconcavity, using the previous equality, there exists $C>0$ s.t.
$0\leq\int_{\mathbb{R}^{3d}}\big<\nabla\psi(x),z_{0}-z_{1}\big>d\alpha_{n}(x,z_{0},z_{1})+C\big(\int_{\mathbb{R}^{2d}}|z_{0}|^{2}d\xi_{n}(x,z_{0})_{+}\int_{\mathbb{R}^{2d}}|z_{1}|^{2}d\widetilde{\xi}_{n}(x,z_{1})\big).$
We then deduce
$0$ $\leq\int_{\mathbb{R}^{3d}}\big<\nabla\psi(x),z_{0}-z_{1}\big>d\alpha_{n}(x,z_{0},z_{1})+C\big(\|\xi_{n}\|^{2}_{\mu}+\|\widetilde{\xi}_{n}\|^{2}_{\mu}\big)$
$\leq\big<(\mathrm{Id}\times\nabla\psi)_{\#}\mu,\xi_{n}\big>_{\mu}-\big<(\mathrm{Id}\times\nabla\psi)_{\#}\mu,\widetilde{\xi}_{n}\big>_{\mu}+C\varepsilon^{2}_{n}$
$=\big<(\mathrm{Id}\times\nabla\psi)_{\#}\mu,\text{Pr}(\xi_{n})\big>_{\mu}-\big<(\mathrm{Id}\times\nabla\psi)_{\#}\mu,\text{Pr}(\widetilde{\xi}_{n})\big>_{\mu}$
Dividing by $\varepsilon_{n}$ and passing to the limit $\varepsilon_{n}\to 0$, we obtain
$\big<(\mathrm{Id}\times\nabla\psi)_{\#}\mu,\xi\big>_{\mu}-\big<(\mathrm{Id}\times\nabla\psi)_{\#}\mu,\widetilde{\xi}\big>_{\mu}\geq 0.$
Consider now the $p\in\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ given at the beginning of the proof: since $W_{\mu}((\mathrm{Id}\times p)_{\mu},(\mathrm{Id}\times\nabla\psi)_{\mu})=\|p-\nabla\psi\|_{L^{2}_{\mu}(\mathbb{R}^{d})}$, a density argument and the alternative description of the tangent space given in (2.11) lead us to
$\big<(\mathrm{Id}\times p)_{\#}\mu,\xi\big>_{\mu}-\big<(\mathrm{Id}\times p)_{\#}\mu,\widetilde{\xi}\big>_{\mu}\geq 0,$
that together with (2.14) gives the desired contraddiction.
The isometric relation (2.13) follows from (2.10), and from the fact that, at absolutely continuous measures, $\text{Pr}(\partial^{-}_{S}\phi(\mu))=\partial_{S,M}^{-}\phi(\mu)$. ∎
2.4. Selection of the differential
In the notion of viscosity solutions and from now on, given $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ proper and lower semicontinuous and $\mu\in\text{Dom}(\phi)$ we adopt the following differential
(2.15) $\partial^{-}\phi(\mu):=\partial^{-}_{S}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}).$
This definition extends to $\partial^{+}$ by the reverse operation $\gamma\mapsto-1\cdot\gamma$, which, as remarked, leaves the tangent space invariant.
Note that the intersection $\partial^{-}\phi(\mu)\cap\partial^{+}\phi(\mu)$ contains at most one element since $\partial^{+}_{S}\phi(\mu)\cap\partial^{-}_{S}\phi(\mu)$ does. Whenever this element exist, we say that the continuous function $\phi$ is differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. This element, denoted with $\partial_{\mu}\phi$, coincides with the Wasserstein gradient (and it is always a map, otherwise $\text{Pr}(\partial_{\mu}\phi)$ would be a different element in the intersection).
We end by stressing that thanks to Lemma (2.12), at absolutely continuous measures, our notion of subdifferential coincides with the one adopted in [DS], up to identification via the barycentric projection, in the definition of viscosity solutions of (1.1). Moreover, being $\partial^{-}\subseteq\partial^{-}_{S,I}$, our subdifferential is a subset of the subdifferential introduced in [BSOT] to treat viscosity solutions of (1.2).
Remark 2.13 (On the extrinsic and intrinsic geometry of the Wasserstein space).
It is worth mentioning that we do not know if $\partial^{-}=\partial^{-}_{K}\cap\mathrm{Tan}^{K}_{(\cdot)}\mathcal{P}_{2}(\mathbb{R}^{d})$, as in the Monge case, see Lemma 2.12. The answer to this question would give a correspondence between the extrinsic (e.g. $\partial^{-}_{S}$) and intrinsic (e.g. $\partial^{-}_{K}$) notions of differential, up to intersect them with the geometric tangent space. We leave this question open for future research.
A similar result which explores the link between extrinsic and intrinsic geometry has been obtained in the recent paper [CavSavSod23], where an equivalence between geodesic convexity (intrinsic convexity) and strong (or total) convexity (extrinsic convexity) has been established for continuous functions, see Theorem 9.1 therein.
We state here an important result (See [BSOT, Proposition 2.4]) that it is fundamental in comparison arguments and describes the superdifferentiability of the Wasserstein distance squared that we need. [^4] Actually, [BSOT, , Proposition 2.4] is stated for $\partial^{+}_{S,I}$. However, a careful inspection of the simple proof reveals that the very same argument gives the result stated here for our choice of superdifferential.
Proposition 2.14.
For any $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ and $\sigma\in\Gamma_{0}(\mu,\nu)$, consider the plan $\xi:=(\pi_{1},\pi_{1}-\pi_{2})_{\#}\sigma$. Then the function $\Phi:\eta\mapsto\frac{1}{2}W^{2}_{2}(\eta,\nu)$ is such that $\xi\in\partial^{+}\Phi(\mu).$
2.5. Main Hypotheses
In what follows, and throughout the paper, we suppose the following
Assumption 2.15.
We assume that the Hamiltonian $H:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ satisfies the following properties:
  • a)
    $H$ is globally continuous;
  • b)
    there exists a constant $C>0$ such that, for all $p,q,x,y\in\mathbb{R}^{d}$ and $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
    $|H(x,p,\mu)-H(y,q,\nu)|\leq C\bigl(1+|p|+|q|\bigr)\bigl(W_{2}(\mu,\nu)+|x-y|+|p-q|\bigr).$
  • c)
    there exists $C>0$
    $|H(x,p,\mu)|\leq C(1+|x|^{2}+|p|^{2}+\mathcal{M}_{2}(\mu)),\quad\forall(x,p,\mu)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
The terminal cost $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is assumed to be bounded and $W_{1}$–Lipschitz continuous.
The diffusion term in (1.2) is of the form $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$
2.6. Definitions of viscosity solutions.
2.6.1. First order HJE
In order to deal with the first-order equation, the author of [BSOT] considers the following relaxed Hamiltonian $\mathcal{H}\colon\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$
(2.16) ${\mathcal{H}}(\gamma)=\int_{\mathbb{R}^{2d}}H(x,z,{\pi_{1}}_{\#}\gamma)d\gamma(x,z).$
This relaxation is reminiscent of the Kantorovich relaxation in the classical optimal transport theory and is the main ingredient of the comparison principle in [BSOT, Theorem 2.11]. We stress that the relaxed Hamiltonian coincides with the usual one at elements induced by maps, and, by the growth condition on $H$,
$\mathcal{H}(\gamma)\leq C(1+\|\gamma\|^{2}_{(\pi_{1})_{\#}\gamma}+\mathcal{M}_{2}((\pi_{1})_{\#}\gamma)),\quad\forall\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}).$
We are ready to give the notion of viscosity solutions for (1.1) by means of the differential defined in (2.15), whose extension to the time dependent case is obvious.
Definition 2.16.
We say that $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}$ is a viscosity subsolution of (1.1) whenever for all $t\in[0,T),\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $(r,\gamma)\in\partial^{+}U(t,\mu)$, we have
$-r+{\mathcal{H}}(\gamma)\leq 0.$
We say that $V:[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}$ is a viscosity supersolution of (1.1) whenever for all $t\in[0,T],\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $(r,\gamma)\in\partial^{-}V(t,\mu)$, we have
$-r+{\mathcal{H}}(\gamma)\geq 0.$
We say that $U$ is a viscosity solution of (1.1) if it is a subsolution and a supersolution.
We say that $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a classical sub(super)solution if $U$ is $C^{1}$ and
$\partial_{t}U(t,\mu)+\int_{\mathbb{R}^{d}}H(x,\mu,\partial_{\mu}U)d\mu\leq(\geq)0,\quad\forall(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
Remark 2.17.
We recall that the notion of viscosity solution introduced in [BSOT, Definition 2.8] is obtained from Definition 2.16 by replacing the differential $\partial$ with the larger set $\partial_{S,I}$. We emphasize that the notion of viscosity solution adopted in Definition 2.16 is weaker than the one considered in [BSOT]. Indeed, any subsolution (resp. supersolution) in the sense of [BSOT, Definition 2.8] is also a subsolution (resp. supersolution) in the sense of Definition 2.16.
On the other hand, existence results for our notion of solution can be obtained more easily, since the definition of viscosity solution only needs to be tested against a smaller class of differentials.
Moreover, the two notions can be compared by means of the comparison principle presented in [BSOT]. In particular, one can prove that they coincide when the underlying space is the flat torus $\mathbb{T}^{d}$, provided that existence holds for both notions of solution. This equivalence follows from Proposition 2.14. For an extension of this result to the whole space $\mathbb{R}^{d}$, one may additionally assume that both solutions are continuous with respect to the weak convergence of measures. This assumption guarantees the existence of a maximum in the doubling variables argument.
2.6.2. Semilinear HJE with idiosyncratic noise
We start by recalling, for the sake of completeness, the definition of Entropy functional and Fisher information. We refer the reader to [AGS] and [OTVillani] for more details.
The entropy functional $\mathcal{E}\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is defined as
$\mathcal{E}(\mu)=\begin{cases}\int_{\mathbb{R}^{d}}\mu\log\mu dx,\quad&\text{if }\mu\in\mathcal{P}_{2,ac}(\mathbb{R}^{d})\\ \infty\quad&\text{otherwise}.\end{cases}$
Here, and throughout the paper, we adopt the usual abuse of notation by denoting by $\mu$ also the density of the measure $\mu$ with respect to the Lebesgue measure $\lambda_{d}$, whenever it exists. Having formally, $\partial_{\mu}\mathcal{E}(\mu)=\nabla\frac{\delta}{\delta\mu}\mathcal{E}(\mu)=\nabla\log\mu$ the Fisher information is defined by
$\mathcal{I}(\mu)=\|\partial_{\mu}\mathcal{E}(\mu)\|_{L^{2}_{\mu}(\mathbb{R}^{d})}^{2}=\int_{\mathbb{R}^{d}}|\nabla\log\mu|^{2}dx=\int_{\mathbb{R}^{d}}\frac{|\nabla\mu|^{2}}{\mu}dx.$
For future use, we recall that
(2.17) $\mathcal{E}^{*}(\mu):=\mathcal{E}(\mu)+\pi\mathcal{M}_{2}(\mu)\geq 0,$
for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, see [DS, Equation  (2.5)].
For the definition of viscosity solution of (1.2), we follow [DS]. Note that the authors introduce an entropic penalization that can be absorbed, thanks to the semilinear structure, into the equation, up to a constant depending only on the magnitude of the penalization itself and the diffusion matrix.
Definition 2.18 (See [DS, Definition 2.10]).
We say that $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}$ is a viscosity subsolution of (1.2) if there is a constant $K>0$, depending only on $U,H$ and the diffusion $a$, such that, for every $\delta>0$, whenever $t\in[0,T],\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathcal{I}(\mu)<\infty$, and $(r,p)\in\partial^{+}\left(U-\delta\mathcal{E}\right)(t,\mu)$, we have
$-r+\int_{\mathbb{R}^{d}}\langle\operatorname{div}_{x}a(\mu,x)+a(\mu,x)\nabla\log\mu(x),p(x)\rangle d\mu(x)+\int_{{\mathbb{R}^{d}}}H\left(x,p(x),\mu\right)d\mu(x)\leq K\delta.$
We say that $V:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}$ is a viscosity supersolution of (1.2) if there is a constant $K>0$, depending only on $V,H$ and the diffusion $a$, such that, for every $\delta>0$, whenever $t\in[0,T],\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathcal{I}(\mu)<\infty$, and $(r,p)\in\partial^{-}\left(V+\delta\mathcal{E}\right)(t,\mu)$, we have
$-r+\int_{\mathbb{R}^{d}}\langle\operatorname{div}_{x}a(\mu,x)+a(\mu,x)\nabla\log\mu(x),p(x)\rangle d\mu(x)+\int_{{\mathbb{R}^{d}}}H\left(x,p(x),\mu\right)d\mu(x)\geq-K\delta.$
We say that $U$ is a viscosity solution of (1.2) if it is both a subsolution and a supersolution.
Remark 2.19.
In view of the computation done in [DS, Section  2.3], the constant $K$ is expected to blow up as the diffusion goes to zero. We will take care of this fact in the proof of the vanishing viscosity limit.
3. Hopf-Lax formula
The goal of this section is to establish an analogue of the finite-dimensional Hopf–Lax formula in the Wasserstein space. We show that this formula provides a representation of the viscosity solution to (1.1) in the sense of Definition 2.16. A first extension in this direction was obtained in [GANGBO2019119] for the quadratic case, where the specific structure of the Hamiltonian is essential due to its relation with a projection operator. Here, we generalize their result beyond the quadratic setting, allowing for a convex Hamiltonian H depending solely on the adjoint variable. Moreover, in contrast with the assumptions in [BSOT, Section 5], we do not require differentiability of the associated Lagrangian. In particular, the Hopf–Lax formula still yields a viscosity solution even when the regularity hypotheses ensuring existence in [BSOT] are not satisfied.
Let us start by recalling the role of the Hopf-Lax formula in finite-dimensional spaces. We consider a Lipschitz continuous datum $g:\mathbb{R}^{d}\to\mathbb{R}$ and a Hamiltonian $H:\mathbb{R}^{d}\to\mathbb{R}$ that depends only on the adjoint variable, convex, superlinear with at most quadratic growth, and has a minimum at 0 (so that its Fenchel transform $L$ is such that $L(0)=0$).
Under these assumptions, it is well known that the function $u:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ defined by the Hopf-Lax formula:
$u(t,x)=\inf_{y\in\mathbb{R}^{d}}\left\{g(y)+(T-t)L\left(\frac{y-x}{T-t}\right)\right\}.$
is the viscosity solution of the HJ equation
$\begin{cases}-\partial_{t}w+H(\nabla w)=0,\quad[0,T)\times\mathbb{R}^{d}\\ w(T,x)=g(x),\quad x\in\mathbb{R}^{d}\end{cases}.$
Note that in the representation formula the case $t=T$, gives $u(T,x)=g(x)$, by the superlinearity of the Lagrangian.
The above formula is formally derived using the method of characteristics. Its power also lies in the fact that it provides an alternative approach to obtain properties such as finite speed propagation of characteristics, Lipschitz continuity of the solution, and the propagation of semiconcavity provided the datum $g(\cdot)$ is semiconcave. Remarkably, in case the Hamiltonian $H$ is strictly convex, the semiconcavity of the solution at positive time is ensured even in the case of a non semiconcave datum. Let us underline that the key ingredient to prove that the function $u$, given by Hopf-Lax formula, is the viscosity solution of the HJ equation is the semigroup structure given in the form of a Dynamic Programming Principle:
(3.1) $u(t,x)=\min_{y\in\mathbb{R}^{d}}\left\{(s-t)L\left(\frac{y-x}{s-t}\right)+u(s,y)\right\},\quad\forall\ 0\leq t<s\leq T.$
The aforementioned properties are standards for finite dimensional Euclidean spaces, see [Evans, Section 3.3], and can be classically proven also for Hilbert spaces, see [lasry_remark_1986] and references therein.
In this section, we make the following assumptions:
  • HopfA
    The function $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is Lipschitz continuous.
  • HopfB
    The Lagrangian $L:\mathbb{R}^{d}\to\mathbb{R}$ is convex, superlinear, and has at most quadratic growth, with $L(0)=0$. Moreover, we assume the following coercivity condition: there exist constants $C,c>0$ such that
    $L(z)\geq c\|z\|^{2}-C\quad\forall\,z\in\mathbb{R}^{d}.$
We start by considering a relaxation of the Lagrangian
3.1. The relaxed Lagrangian as the Fenchel transform of the relaxed Hamiltonian
$\mathcal{L}\colon\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$, defined by
(3.2) $\mathcal{L}(\gamma):=\int_{\mathbb{R}^{2d}}L\left(z\right)\gamma(dx,dz).$
We observe that the following property holds
(3.3) $\mathcal{L}(\gamma)\geq c\|\gamma\|^{2}_{(\pi_{1})_{\#}\gamma}-C,$
for $C,c>0$ the same constant of the coercivity condition of the Lagrangian $L$.
The link between the relaxed Lagrangian and the relaxed Hamiltonian is explained by the following
Lemma 3.1 (Fenchel Transform of the Relaxed Hamiltonian).
Let $\mathcal{H}:\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$ be the relaxed Hamiltonian associated to $H(p)=\sup_{v\in\mathbb{R}^{d}}\big\{\big<-v,p\big>-L(v)\big\}$, i.e. $L^{*}=H$ defined in (2.16). Fix $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then, the equalities
$\mathcal{H}(\gamma)$ $=\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v)\Big\}$
$=\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\big<-\xi,\gamma\big>_{\theta,\mu}-\mathcal{L}(\xi)\Big\}$
$=\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\Big\{\big<-\xi,\gamma\big>_{\mu}-\mathcal{L}(\xi)\Big\}$
hold for every choice of $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$.
Moreover, for $\theta\in\Gamma_{\mu}(\gamma,\xi)$
(3.4) $\mathcal{H}(\gamma)+\mathcal{L}(\xi)=\big<-\xi,\gamma\big>_{\theta,\mu}\Longleftrightarrow\mathrm{supp}(\theta)\subseteq\big\{(x,p,v)\in\mathbb{R}^{3d}:p\in-D^{-}L(v)\big\},$
where $D^{-}L$ denotes the subdifferential of $L$ in the classic convex sense, and the set of $\theta\in\Gamma_{\mu}(\gamma,\xi)$ satisfying (3.4) is not empty since it contains at least an element induced by a measurable map.
Proof.
Let us prove the first equality, the second and the third one being immediate from this one.
Fix $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$. By the very definition of relaxed Hamiltonian (2.16) and Fenchel transform we have
$\mathcal{H}(\gamma)=\int_{\mathbb{R}^{2d}}H(p)d\gamma(x,p)$ $\underbrace{=}_{\theta\in\Gamma_{\mu}(\gamma,\xi)}\int_{\mathbb{R}^{3d}}H(p)d\theta(x,p,v)$
$\underbrace{\geq}_{L^{*}=H}\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v),\quad\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\,\theta\in\Gamma_{\mu}(\gamma,\xi).$
Equivalently, $\mathcal{H}(\gamma)\geq\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v)\Big\}$. The equality holds iff $\theta$ is concentrated on the multigraph $\text{Graph}({-D^{-}L})=\big\{(x,p,v)\in\mathbb{R}^{3d}:p\in-D^{-}L(v)\big\}$ by the characterization of optimality in the Fenchel transform. Note that $\text{Graph}({-D^{-}L})$ is closed, and it is not empty, since $L$ is convex and superlinear. Thereofore, for a fixed $\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d})$, we can find a Borel measurable function $f:\mathbb{R}^{2d}\to\mathbb{R}^{d}$ s.t.
$(x,f(x,v),v)\in\text{Graph}({-D^{-}L})\quad\gamma-(x,v)\,a.e.$
see Kuratowski-Ryll-Nardzewski Theorem [Bogachev2007, Thm 6.9.3]. In particular, $\theta:=(\text{Id}\times f\times\text{Id})_{\#}{\gamma}$ is s.t. $(\pi_{1},\pi_{2})_{\#}\theta=\gamma$ and it is concentrated on $\text{Graph}({-D^{-}L})$.
3.2. Hopf-Lax Formula, Dynamic Programming Principle, regularity and infinitesimal behavior of optimizers
For the sake of presentation, we derive the Hopf-Lax formula from the associated control problem. Consider the value function considered in [BSOT]
$U_{K}(t,\mu)=\inf_{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)}{\int_{t}^{T}\mathcal{L}(\zeta_{s})ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})},$
where for $t<T$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
$\mathcal{A}_{[t,T]}(\mu):=\Big\{\zeta_{\cdot}\in C([t,T];\mathcal{P}_{2}(\mathbb{R}^{2d})):\partial_{s}\mu_{s}+\text{div}(\text{Pr}(\zeta_{s})\mu_{s})=0,\,\mu_{s}=(\pi_{1})_{\#}\zeta_{s},\,\mu_{t}=\mu\,\,\,\forall s\in[t,T]\Big\}.$
Let us show that this value function coincides with the classical value function $U_{M}$ defined as
$U_{M}(t,\mu):=\inf_{(\mu_{\cdot},\alpha_{\cdot})\in\mathcal{A}^{M}_{[t,T]}(\mu)}{\int_{t}^{T}\int_{\mathbb{R}^{d}}L(\alpha_{s}(x))d\mu_{s}ds+\mathcal{G}(\mu_{T})},$
[…]
$\mathcal{A}^{M}_{[t,T]}(\mu):=\Big\{(\mu_{\cdot},\alpha_{\cdot})\in C([t,T];\mathcal{P}_{2}(\mathbb{R}^{2d})\times L^{2}_{\mu_{\cdot}}(\mathbb{R}^{d};\mathbb{R}^{d})):\partial_{s}\mu_{s}+\text{div}(\alpha_{s}\mu_{s})=0,\,\,\,s\in[t,T];\,\mu_{t}=\mu,\Big\}.$
Indeed, we note that
$U_{K}(t,\mu)\leq\inf_{\big\{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)\,:\,\zeta_{s}=\mu_{s}\otimes\delta_{\alpha_{s}}\big\}}{\int_{t}^{T}\mathcal{L}(\zeta_{s})ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})}=U_{M}(t,\mu).$
On the other hand, by Jensen inequality $\mathcal{L}(\zeta)\geq\mathcal{L}((\mathrm{Id}\times\Pr(\zeta))),\,\forall\zeta\in\mathcal{P}_{2}(\mathbb{R}^{2d})$. Therefore, $\forall(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$
$U_{K}(t,\mu)\geq\inf_{\zeta_{\cdot}\in\mathcal{A}_{[t,T]}(\mu)}{\int_{t}^{T}\mathcal{L}(\mathrm{Id}\times\Pr(\zeta_{s}))ds+\mathcal{G}((\pi_{1})_{\#}\zeta_{T})}\geq U_{M}(t,\mu).$
Therefore, $U_{M}=U_{K}$.
In view of the previous computations, the finite dimensional Hopf-Lax formula (3.1) can naturally be generalized in the following way.
Definition 3.2 (Hopf-Lax formula).
Fix $\mathcal{G}\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$, l.s.c. and bounded from below. We define the function
$\mathcal{V}(t,\mu):$ $=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{d})}\left\{\mathcal{G}((\pi_{2})_{\#}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot(\pi_{1},\pi_{2}-\pi_{1})_{\#}\gamma\right)\right\}$
where $t\in[0,T],\,\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, as the Hopf-Lax evolution of $\mathcal{G}$.
Remark 3.3.
Firslty, we observe that (3.2) can be equivalently rewritten as
$\mathcal{V}(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{d})}\left\{\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\right\}.$
Then, as in finite dimension, we also remark that thanks to the coercivity property (3.3), when $t=T$, we have
(3.6) $\lim_{t\to T^{-}}(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)=\begin{cases}0\quad&\text{if }\gamma=\mu\otimes\delta_{o}\\ +\infty\quad&\text{otherwise}\end{cases}.$
Therefore, $\mathcal{V}(T,\mu)=\mathcal{G}(\exp(\mu\otimes\delta_{o}))=\mathcal{G}(\mu)$.
Moreover, since $\mathcal{G}$ is bounded from below, $\mathcal{V}>$-$\infty$. It is also immediate to see that the minimization can be split to obtain
$\mathcal{V}(t,\mu)=\inf_{\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{\mathcal{G}(\nu)+(T-t)\mathcal{L}_{T-t}(\mu,\nu)\right\},$
where $\mathcal{L}_{T-t}(\mu,\nu):=\min_{\gamma\in\Gamma(\mu,\nu)}\Big\{\int_{\mathbb{R}^{2d}}L\left(\frac{y-x}{T-t}\right)d\gamma(x,y)\Big\}.$
Let us now prove that the function $\mathcal{V}$ defined in (3.2) satisfies a semigroup property.
Proposition 3.4 (Dynamic Programming Principle).
Let $L:\mathbb{R}^{d}\to\mathbb{R}$ satisfy (HopfB). Then
(3.7) $\mathcal{V}(t,\mu)=$ $\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{V}(s,\exp_{\mu}(\gamma))+(s-t)\mathcal{L}\left(\frac{1}{s-t}\cdot\gamma\right)\right\},\quad\forall\,0\leq t<s<T.$
Proof.
Denote by $\mathcal{K}(t,\mu)$ the RHS of (3.7). In this proof, we use the subscript in a different way: $\gamma_{t}:=(\pi_{1},\frac{\pi_{2}-\pi_{1}}{t})_{\#}\gamma,$ where $t>0$ and $\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d})$. Then $\mathcal{K}(t,\mu)$ coincide with
$\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{V}(s,(\pi_{2})_{\#}\gamma)+(s-t)\mathcal{L}(\gamma_{s-t})\right\},\quad\forall\,0\leq t<s<T.$
We then focus on proving the equality
$\mathcal{V}(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{V}(s,(\pi_{2})_{\#}\gamma)+(s-t)\mathcal{L}(\gamma_{s-t})\right\}\quad\forall\,0\leq t<s<T.$
Fix $0\leq t<T$ and $\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$. For $t<s<T$ and $\gamma^{1\,2}\in\mathcal{P}_{2,\mu_{1}}(\mathbb{R}^{2d})$, let us denote $\mu_{2}:=(\pi_{2})_{\#}\gamma^{1\,2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and use the formula of $\mathcal{V}$ given in Definition 3.2, so that for $\varepsilon>0$, there exists $\gamma^{2\,3}\in\Gamma(\mu_{2},\mu_{3})$ such that
(3.8) $\mathcal{V}(s,\mu_{2})\geq(T-s)\mathcal{L}\left(\gamma^{2\,3}_{T-s}\right)+\mathcal{G}(\mu_{3})-\varepsilon,$
while, for every $\gamma\in\Gamma(\mu_{1},\mu_{3})$ it holds
(3.9) $\mathcal{V}(t,\mu_{1})\leq(T-t)\mathcal{L}(\gamma_{T-t})+\mathcal{G}(\mu_{3}).$
Since $T>s>t$, we have that $\frac{T-s}{T-t}+\frac{s-t}{T-t}=1$ with addenda in $(0,1)$. Therefore, by the convexity of $L$, we obtain
(3.10) $L\left(\frac{z-x}{T-t}\right)\leq\frac{s-t}{T-t}L\left(\frac{y-x}{s-t}\right)+\frac{T-s}{T-t}L\left(\frac{z-y}{T-s}\right).$
Now, let $\gamma^{1\,3}\in\Gamma(\mu_{1},\mu_{3})$ be the composition of plans defined in [[AGS], Remark 5.3.3] as the plan given by
$\gamma^{1\,3}:=\gamma^{2\,3}\circ\gamma^{1\,2}(dx,dz)=\int_{\mathbb{R}^{d}}\left(\gamma_{y}^{2\,3}(dz)\times\gamma_{y}^{1\,2}(dx)\right)\mu_{2}(dy)$
where $\gamma^{1\,2}(dx,dy)=\gamma_{y}^{1\,2}(dx)\mu_{2}(dy)$ and $\gamma^{2\,3}(dy,dz)=\gamma_{y}^{2\,3}(dz)\mu_{2}(dy)$. Then we can integrate (3.10) with respect to $\gamma^{1\,3}$ to obtain the following “triangular Inequality”
(3.11) $\mathcal{L}\left(\gamma^{1\,3}_{T-t}\right)\leq\frac{s-t}{T-t}\mathcal{L}\left(\gamma^{1\,2}_{s-t}\right)+\frac{T-s}{T-t}\mathcal{L}\left(\gamma^{2\,3}_{T-s}\right).$
Thus combining (3.9) with (3.11) and using (3.8), we have
$\mathcal{V}(t,\mu_{1})$ $\leq(s-t)\mathcal{L}\left(\gamma^{1\,2}_{s-t}\right)+(T-s)\mathcal{L}\left(\gamma^{2\,3}_{T-s}\right)+\mathcal{G}(\mu_{3})$
$\leq(s-t)\mathcal{L}\left(\gamma^{1\,2}_{s-t}\right)+\mathcal{V}(s,\mu_{2})+\varepsilon$
By the arbitrariness of $\varepsilon$ we obtain $\mathcal{V}(t,\mu_{1})\leq\mathcal{K}(t,\mu_{1})$.
For the converse inequality we proceed as follows: fix $0\leq t<T$ and $\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $\mu_{3}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\gamma^{1\,3}\in\Gamma(\mu_{1},\mu_{3})$ be $\varepsilon$-optimizers for $\mathcal{V}(t,\mu_{1})$.
For a fixed $t<s<T$, consider
$\mu_{2}:=\left(\frac{T-s}{T-t}\pi_{1}+\left(1-\frac{T-s}{T-t}\right)\pi_{2}\right)_{\#}\gamma^{1\,3},$
then, for
$\gamma^{2\,3}:=\left(\frac{T-s}{T-t}\pi_{1}+\left(1-\frac{T-s}{T-t}\right)\pi_{2},\pi_{2}\right)_{\#}\gamma^{1\,3}\in\Gamma(\mu_{2},\mu_{3})$
and
$\gamma^{1\,2}=:\left(\pi_{1},\frac{T-s}{T-t}\pi_{1}+\left(1-\frac{T-s}{T-t}\right)\pi_{2}\right)_{\#}\gamma^{1\,3}\in\Gamma(\mu_{1},\mu_{2}),$
we have
$\mathcal{K}(t,\mu_{1})\leq(s-t)\mathcal{L}\left(\gamma^{1\,2}_{s-t}\right)+\mathcal{V}(s,\mu_{2})$ $\leq(s-t)\mathcal{L}\left(\gamma^{1\,2}_{s-t}\right)+(T-s)\mathcal{L}\left(\gamma^{2\,3}_{T-s}\right)+\mathcal{G}(\mu_{3})$
$=(T-t)\mathcal{L}\left(\gamma^{1\,3}_{T-t}\right)+\mathcal{G}(\mu_{3})\leq\mathcal{V}(t,\mu_{1})+\varepsilon,$
where we used the fact that $\mathcal{L}\left(\gamma^{1\,3}_{T-t}\right)=\mathcal{L}\left(\gamma^{2\,3}_{T-s}\right)=\mathcal{L}\left(\gamma^{1\,2}_{s-t}\right)$.
Minimizing the LHS of the previous equation gives the desired result. ∎
Lemma 3.5.
The function $\mathcal{V}$ is Lipschitz continuous in $0,\infty)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, provided that ${\mathrm{Lip}}(\mathcal{G},W_{2})<\infty$. More precisely, for all $0<t<T$
$\mathrm{Lip}(\mathcal{V}(t,\cdot);W_{2})\leq\mathrm{Lip}(\mathcal{G};W_{2})$
and for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$
$\mathrm{Lip}_{t}(\mathcal{V}(\cdot,\mu))\leq\sup\Big\{\mathcal{H}(\xi)\,|\,\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\|\xi\|_{\mu}\leq\mathrm{Lip}(\mathcal{G};W_{2})\Big\}.$
Proof.
The proof is an adaptation of the proof of [Evans, Lemma 2, Chapter 3] to the Wasserstein framework.
Fix $0<t<T,\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Fix $\varepsilon>0$, and choose $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ s.t.
$\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\leq\mathcal{V}(t,\mu)+\varepsilon,$
and pick $\theta\in\Gamma_{0}(\gamma,\nu).$ We note that $\xi:=(\pi_{3},\pi_{2}+\pi_{3}-\pi_{1})_{\#}\theta\in\mathcal{P}_{2,\nu}(\mathbb{R}^{d})$, and $\mathcal{L}(\zeta_{T-t})=\mathcal{L}(\frac{1}{T-t}\cdot\gamma)$, where for $\zeta_{T-t}$ we adopted the same notation used in the proof of the DPP 3.4. Therefore,
$\mathcal{V}(t,\nu)-\mathcal{V}(t,\mu)$ $\leq\inf_{\zeta\in\mathcal{P}_{2,\nu}(\mathbb{R}^{2d})}\Big\{\mathcal{G}((\pi_{2})_{\#}\zeta)+(T-t)\mathcal{L}\left(\zeta_{T-t}\right)\Big\}-\mathcal{G}(\exp_{\mu}(\gamma))-(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)+\varepsilon$
${\underbrace{\leq}_{\zeta=\xi}}\mathcal{G}((\pi_{2})_{\#}\zeta)-\mathcal{G}(\exp_{\mu}(\gamma))+\varepsilon\leq\mathrm{Lip}(\mathcal{G};W_{2})W_{2}((\pi_{2})_{\#}\zeta,\exp_{\mu}(\gamma))+\varepsilon$
$=\mathrm{Lip}(\mathcal{G};W_{2})(\int_{\mathbb{R}^{3d}}|z-x|^{2}d\theta(x,y,z))^{\frac{1}{2}}+\varepsilon$
$=\mathrm{Lip}(\mathcal{G};W_{2})W_{2}(\mu,\nu)+\varepsilon.$
Since $\varepsilon$ is arbitrary, we can pass to the limit by letting $\varepsilon\to 0$, then inverting the role of the measures $\mu$ and $\nu$, we get the desired bound. In particular, $\mathrm{Lip}(\mathcal{V}(t,\cdot);W_{2})$ is time independent.
We now prove the time Lipschitz estimate. By the DPP, Proposition 3.4, fixing $0<t<t+h<T$ for $h>0$ and choosing $\gamma=(\mu\otimes\delta_{o})\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ we have
(3.12) $\mathcal{V}(t,\mu)\leq\mathcal{V}(t+h,\mu)+h\mathcal{L}\left(\frac{1}{h}\cdot(\mu\otimes\delta_{o})\right)=\mathcal{V}(t+h,\mu),$
where we used the fact that $L(0)=0$.
On the other hand, using again the DPP and $\mathrm{Lip}(\mathcal{V}(t,\cdot);W_{2})\leq\mathrm{Lip}(\mathcal{G};W_{2})$, we have
$\mathcal{V}(t,\mu)$ $=\inf_{\zeta\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{V}(t+h,\exp_{\mu}(\zeta))+h\mathcal{L}\left(\frac{1}{h}\cdot\zeta\right)\right\}$
$\geq\inf_{\zeta\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{V}(t+h,\mu)+h\mathcal{L}\left(\frac{1}{h}\cdot\zeta\right)-\mathrm{Lip}(\mathcal{G};W_{2})W_{2}(\mu,\exp_{\mu}(\zeta))\right\}$
(3.13) $=\mathcal{V}(t+h,\mu)+\inf_{\zeta\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{h\mathcal{L}\left(\frac{1}{h}\cdot\zeta\right)-\mathrm{Lip}(\mathcal{G};W_{2})W_{2}(\mu,\exp_{\mu}(\zeta))\right\}.$
In particular, we have by the one sided Cauchy-Schartz inequality
$\mathrm{Lip}(\mathcal{G};W_{2})W_{2}(\mu,\exp_{\mu}(\zeta))$ $\leq\mathrm{Lip}(\mathcal{G};W_{2})\|\zeta\|_{\mu}$
$=\max\Big\{\big<\xi,\zeta\big>_{\mu}:{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\|\xi\|_{\mu}\leq\mathrm{Lip}(\mathcal{G};W_{2})}\Big\}$
(3.14) $=\max\Big\{\big<-\xi,\zeta\big>_{\mu}:{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\|\xi\|_{\mu}\leq\mathrm{Lip}(\mathcal{G};W_{2})}\Big\}.$
The combination of (3.13) and (3.14) gives,
$\mathcal{V}(t,\mu)$ $\geq\mathcal{V}(t+h,\mu)+\inf_{\zeta\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{h\mathcal{L}\left(\frac{1}{h}\cdot\zeta\right)-\sup\Big\{\big<-\xi,\zeta\big>_{\mu}:{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\|\xi\|_{\mu}\leq\mathrm{Lip}(\mathcal{G};W_{2})}\Big\}\right\}$
$=\mathcal{V}(t+h,\mu)-h\sup\left\{-\mathcal{L}\left(\frac{1}{h}\cdot\zeta\right)+\big<-\xi,\frac{1}{h}\cdot\zeta\big>_{\mu}\,\big|\,\zeta,\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\|\xi\|_{\mu}\leq\mathrm{Lip}(\mathcal{G};W_{2})\right\}$
$=\mathcal{V}(t+h,\mu)-h\sup\Big\{\mathcal{H}(\xi)\,|\,\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\|\xi\|_{\mu}\leq\mathrm{Lip}(\mathcal{G};W_{2})\Big\}.$
Summarizing,
[…]
where the RHS is finite thanks to the quadratic growth assumption on the Hamiltonian. ∎
Lemma 3.6 (Infinitesimal behavior of optimizers).
Let $0<t<t+h<T$ for $h>0$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Fix $\varepsilon>0$, and let $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ be an $\varepsilon$-optimizer of
$\mathcal{V}(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{V}(t+h,\exp_{\mu}(\gamma))+h\mathcal{L}\left(\frac{1}{h}\cdot\gamma\right)\right\}.$
If $\text{Lip}(\mathcal{G};W_{2})<\infty$, then there exists a constant $C>0$, depending only on $\text{Lip}(\mathcal{G};W_{2})$ and on the coercitivity parameters of $L$, s.t.
$\|\xi\|_{\mu}\leq C(h+\sqrt{\varepsilon h}).$
Proof.
Let $0<t<t+h<T$ for $h>0$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Fix $\varepsilon>0$ and let $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ be an $\varepsilon$-optimizer. We omit here the dependence of $\xi$ on $h$. By the minimality property and (3.12), we deduce
$\mathcal{V}(t+h,\exp_{\mu}(\xi))+h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right)\leq\mathcal{V}(t,\mu)+\varepsilon\leq\mathcal{V}(t+h,\mu)+\varepsilon,$
hence
$0\leq\mathcal{V}(t+h,\mu)-\mathcal{V}(t+h,\exp_{\mu}(\xi))-h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right)+\varepsilon.$
By the Lipschitz property in Lemma 3.5, we have
[…]
Moreover, by the coercitivity assumption on the Lagrangian,
$Ch+c\frac{1}{h}\|\xi\|^{2}_{\mu}\leq h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right)\leq\mathrm{Lip}(\mathcal{G};W_{2})W_{2}(\mu,\exp_{\mu}(\xi))+\varepsilon\leq\mathrm{Lip}(\mathcal{G};W_{2})\|\xi\|_{\mu}+\varepsilon.$
By a standard absorbing argument we deduce
$\frac{1}{h}\|\xi\|^{2}_{\mu}\leq C(h+\varepsilon),$
with $C$ depending on $\mathcal{G}$ and on the growth parameters of $L$, and using the subadditivity of the square root, we conclude the claim.
3.3. Existence result
We show now that the Hopf-Lax formula gives a viscosity solution in the sense of Definition 2.16 for equation (1.1).
3.3.1. Subsolution property
Fix $(t,\mu)\in(0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $(r,\gamma)\in\partial^{+}U(t,\mu)$, namely $\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
$U(t+h,\exp_{\mu}(h\cdot\xi))-U(t,\mu)\leq h\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(h\|\xi\|_{\mu})+o(h),\quad\forall h>0\text{ small enough},$
where $\theta\in\Gamma_{\mu}(\gamma,\xi).$ By the DPP, we have
$U(t,\mu)\leq U(t+h,\exp_{\mu}(\widetilde{\xi}))+h\mathcal{L}\left(\frac{1}{h}\cdot\widetilde{\xi}\right)\quad\forall\widetilde{\xi}\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}).$
In particular, by choosing $\widetilde{\xi}=h\cdot\xi$ and combining the two previous inequalities, we get
$0\leq h\mathcal{L}(\xi)+h\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(h)+o(h\|\xi\|_{\mu}),\quad\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),h>0\text{ small enough}.$
Dividing by $h>0$ and sending $h\to 0$, we get
$0\leq\mathcal{L}(\xi)+\big<\gamma,\xi\big>_{\theta,\mu}+r.$
Therefore, since $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ is arbitrary, we can invoke Lemma 3.1 to obtain
$0\leq-\mathcal{H}(\gamma)+r,$
the desired inequality.
3.3.2. Supersolution Property
Fix $(t,\mu)\in(0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $(r,\gamma)\in\partial^{-}U(t,\mu)$. Therefore, $\forall\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ we have
$U(t+h,\exp_{\mu}(\xi))-U(t,\mu)\geq\big<\gamma,\xi\big>_{\theta,\mu}+rh+o(\|\xi\|_{\mu})+o(h),\quad\forall h>0\text{ small enough.}$
For a fixed $h>0$, let $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ be an $\varepsilon$-optimizer of (3.7), with $\varepsilon=h$ i.e.
$U(t,\mu)+h\geq U(t+h,\exp_{\mu}(\xi))+h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right).$
We do not stress here, as before, the dependence of $\xi$ on $h>0$. In any case, we know by Lemma 3.6 that $\|\xi\|_{\mu}\leq 2Ch$, with $C$ specified therein. Therefore $o(\|\xi\|_{\mu})=o(h)$.
Combining the two inequalities above at such a $\xi$ , we obtain
$h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right)\geq h\big<\gamma,\frac{1}{h}\cdot\xi\big>_{\theta,\mu}+rh+o(h).$
By the Fenchel inequality in Lemma 3.1, we have
$h\mathcal{L}\left(\frac{1}{h}\cdot\xi\right)+h\big<\gamma,-\frac{1}{h}\cdot\xi\big>_{\theta,\mu}\leq h\mathcal{H}(\gamma),$
hence we obtain
$rh+h\mathcal{H}(\gamma)\geq o(h).$
We conclude by dividing by $h>0$ and passing to the limit as $h\to 0$.
4. Useful Estimates and Lemmata
In this section, we derive some results that will be used in the main argument for the proof of the vanishing viscosity limit.
We first start by showing that, as a consequence of the Benamou-Brenier control representation, the elements of the subdifferential of $W_{2}$-Lipschitz functions are bounded:
Lemma 4.1.
($L^{2}$ version of [DS, Lemma  3.4]) Let $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a $W_{2}$- Lipschitz function and set $L=\mathrm{Lip}(U;W_{2})<\infty$. Let $\mu_{0}\in\mathcal{P}_{2,ac}(\mathbb{R}^{d})$ be such that $\mathcal{E}(\mu_{0})<\infty$ and let $\delta\geq 0$, and $p\in\partial^{+}(U-\delta\mathcal{E})(\mu_{0})$ (resp. $p\in\partial^{-}(U+\delta\mathcal{E})(\mu_{0})$), then $\|p+\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$ (resp. $\|p-\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$)
Proof.
Since the proof is very similar to the one given in [DS, Lemma  3.4], we only give a sketch of it in the superdifferential case.
Let $v:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be smooth and bounded, and consider, for some $t_{0}\geq 0$, the flow given by the continuity equation
(4.1) $\partial_{t}\mu_{t}+\operatorname{div}(v(x)\mu_{t})=0\quad\text{in }(t_{0},\infty),\quad\mu|_{t=t_{0}}=\mu_{0}.$
In view of the smoothness of $v$, the solution $t_{0},T]\ni t\mapsto\mu_{t}\in L^{1}({\mathbb{R}^{d}})$ is continuous.
Moreover, by the representation theorem for solutions of the continuity equation given in [AGS, Thm. 8.3.1] we have
[…]
where $|\mu^{\prime}|(s)$ denotes the metric derivative w.r.t. time in $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ at time $s\in(t_{0},t)$ of the absolutely continuous curve $t\mapsto\mu_{t}$. The previous bound and $L=\mathrm{Lip}(U;W_{2})<\infty$, give
$|U(\mu_{t})-U(\mu_{0})|\leq LW_{2}(\mu_{t},\mu_{0})\leq L\int_{t_{0}}^{t}\|v\|_{L^{2}_{\mu_{s}}(\mathbb{R}^{d})}ds.$
Moreover, by the renormalization property satisfied by $v$ and standard regularization and localization procedures (see [DS, Lemma  3.4]), we also have
$\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{0})=-\int_{t_{0}}^{t}\int_{\mathbb{R}^{d}}\text{div}\,v(x)d\mu_{s}ds.$
Fix now $h,\varepsilon>0$ and $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. By similar estimates performed in [DS, Lemma  3.4], it follows, from the superdifferential property and the Cauchy-Schwarz inequality, that for all sufficiently small $h>0$, depending on $\|v\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}$ and $\varepsilon$,
$\int_{t_{0}}^{t_{0}+h}\int_{\mathbb{R}^{d}}\left(\langle\nabla\varphi(x),v(x)\rangle-\delta\operatorname{div}v(x)\right)d\mu_{s}(x)$
$\geq-L\int_{t_{0}}^{t_{0}+h}\|{v}\|_{L^{2}_{\mu_{s}}(\mathbb{R}^{d})}ds-\|\nabla^{2}\varphi\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}o(h)-\left\|\nabla\varphi-p\right\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\int_{t_{0}}^{t_{0}+h}\|{v}\|_{L^{2}_{\mu_{s}}(\mathbb{R}^{d})}ds-\varepsilon h,$
where $p\in\partial^{+}(U-\delta\mathcal{E})(\mu_{0})$ has been taken as in the statement.
Dividing by $h$ and sending $h\to 0^{+}$, due to the fact that $v$ and $\varphi$ are smooth and bounded, and $\mu_{s}\to\mu_{0}$ as $s\to t_{0}^{+}$ in $L^{1}({\mathbb{R}^{d}})$, we obtain
$\int_{\mathbb{R}^{d}}\left(\langle\nabla\varphi(x),v(x)\rangle-\delta\operatorname{div}v(x)\right)d\mu_{0}(x)\geq-L\|v\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}-\left\|\nabla\varphi-p\right\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\|{v}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}-\varepsilon.$
Since $\varepsilon$ and $\varphi$ were arbitrary and $p\in\mathrm{Tan}^{M}_{\mu_{0}}\mathcal{P}_{2}({\mathbb{R}^{d}})$, we infer that
$\int_{\mathbb{R}^{d}}\left(\big<p(x),v(x)\big>-\delta\operatorname{div}v(x)\right)d\mu_{0}(x)\geq-L\|v\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}.$
In other words, by the arbitrariness of $v$, we conclude that the distribution
$p+\delta\nabla\log\mu_{0}=p+\delta\frac{\nabla\mu_{0}}{\mu_{0}}$
is bounded in $L^{2}_{\mu_{0}}(\mathbb{R}^{d})$ with operator norm $L$. By the $L_{\mu_{0}}^{2}(\mathbb{R}^{d})$ duality, this element belongs to $L^{2}_{\mu_{0}}(\mathbb{R}^{d})$ and $\|p+\delta\nabla\log\mu_{0}\|_{L^{2}_{\mu_{0}}(\mathbb{R}^{d})}\leq L$. ∎
The following Lemma will be directly applied in the proof of the vanishing viscosity theorem.
Lemma 4.2 (Analogue of Proposition 3.1 of [DS] in our context.).
Let $U,V:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be two bounded functions s.t. $\mathrm{Lip}(U;W_{1})+\mathrm{Lip}(V;W_{2})<\infty$ and suppose $V$ is l.s.c. with respect to the $W_{p}$ topology for some $p\in[1,2)$. Fix $\alpha,\delta>0$, then the function $\Phi:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[-\infty,+\infty)$ defined by
$\Phi(\mu,\nu)=U(\mu)-V(\nu)-\frac{1}{2\alpha}W_{2}(\mu,\nu)-\delta\mathcal{E}(\mu)-\delta\pi\mathcal{M}_{2}(\mu)-\delta\pi\mathcal{M}_{2}(\nu)$
achieves its supremum. Moreover, any maximum point $(\overline{\mu},\overline{\nu})$ belongs to $\mathcal{P}_{2,ac}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$, and
  • Set
    $p(x):=\frac{1}{\alpha}(x-T_{\overline{\mu}}^{\overline{\nu}}(x))+2\delta\pi x,\quad x\in\mathbb{R}^{d}$
    $\xi:=(\pi_{1},\frac{\pi_{1}-\pi_{2}}{\alpha}-(2\delta\pi)\pi_{1})_{\#}\sigma\in\mathrm{Tan}^{K}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$
    where $\sigma\in\Gamma_{0}(\overline{\nu},\overline{\mu})$ is such that $(\pi_{2},\pi_{1})_{\#}\sigma=(\mathrm{Id}\times T_{\overline{\mu}}^{\overline{\nu}})_{\#}\overline{\mu}\in\Gamma_{0}(\overline{\mu},\overline{\nu}).$
    Then $p\in{\partial^{+}}\big(U-\delta\mathcal{E}\big)(\overline{\mu})$ and $\xi\in\partial^{-}V(\overline{\nu})$. Moreover,
    $\|p+\delta\nabla\log\overline{\mu}\|_{L^{\infty}}\leq\mathrm{Lip}(U;W_{1}),\,\|p+\delta\nabla\log\overline{\mu}\|_{L^{2}_{\overline{\mu}}(\mathbb{R}^{d})}\leq\mathrm{Lip}(U;W_{2})\text{ and }\|\xi\|_{\overline{\nu}}\leq\mathrm{Lip}(V;W_{2}).$
  • $\overline{\mu},\overline{\mu}^{-1}\in W^{1,\infty}_{\mathrm{loc}}(\mathbb{R}^{d})$ and $\mathcal{I}(\overline{\mu})<\infty$.
Proof.
The attainment of the maximum follows the same argument used in the proof of [DS, Proposition 3.1]. The moment penalization forces any maximizing sequence to remain in a compact (for the weak convergence) region of $\mathcal{P}_{2}(\mathbb{R}^{d})$, so that it admits a weakly convergent subsequence. By the assumed regularity of $U$ and $V$, and recalling the l.s.c. of the entropy w.r.t. the weak convergence, the functional $\Phi$ is upper semicontinuous with respect to weak convergence. Therefore, the limit of such a subsequence is indeed a maximizer.
Let $(\overline{\mu},\overline{\nu})\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$ be a maximizer for $\Phi$. We note that $p\in\partial^{+}(U-\delta\mathcal{E})(\overline{\mu})$ has been already proved in [DS, Proposition 3.1].
The fact that $\xi$ belongs to the subdifferential $\partial^{-}V(\overline{\nu})$ follows directly from Proposition 2.14. Moreover, since $2\pi\delta\nabla|x|^{2}\in\mathrm{Tan}^{M}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$, we also have $\xi\in\mathrm{Tan}^{K}_{\overline{\nu}}\mathcal{P}_{2}(\mathbb{R}^{d})$. Finally, observe that
$\|\gamma\|_{\overline{\nu}}\,\mathrm{Lip}(V;W_{2})\geq W_{2}(\overline{\nu},\exp_{\overline{\nu}}(\gamma))\,\mathrm{Lip}(V;W_{2})\geq V(\exp_{\overline{\nu}}(\gamma))-V(\overline{\nu})\geq\langle\xi,\gamma\rangle_{\theta,\overline{\nu}}+o(\|\gamma\|_{\overline{\nu}}),$
which yields the claim by choosing $\gamma=\frac{\varepsilon}{\|\xi\|_{\overline{\nu}}}\,\xi$ with $\varepsilon>0$, dividing by $\varepsilon$ and letting $\varepsilon\to 0$.
We now focus on the regularity of the optimizer $\overline{\mu}$. The proof is similar to that done in [DS, Propositon  3.7], and we just summarize the argument. Since $\mathrm{Lip}(U;W_{1})<\infty$, [DS, Lemma 3.3] yields $\overline{\mu}^{-1}\in L^{\infty}_{\text{loc}}.$ Therefore, $T_{\overline{\mu}}^{\overline{\nu}}$ is locally bounded, being a.e.-$\overline{\mu}$ the gradient of an everywhere defined convex function. In particular, $p$ is locally bounded. So that, we have
$\begin{cases}p\in L^{2}_{\overline{\mu}}(\mathbb{R}^{d})\\ p+\delta\nabla\log\overline{\mu}\in L^{2}_{\overline{\mu}}(\mathbb{R}^{d}),\quad\end{cases}\Longrightarrow\nabla\log{\overline{\mu}}\in L^{2}_{\overline{\mu}}(\mathbb{R}^{d}),\text{ i.e. }\mathcal{I}(\overline{\mu})<\infty$
$\begin{cases}p\in L_{\mathrm{loc}}^{\infty}(\mathbb{R}^{d})\\ p+\delta\nabla\log\overline{\mu}\in L^{\infty}(\mathbb{R}^{d})\end{cases}\Longrightarrow\nabla\log{\overline{\mu}}\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{d})\underbrace{\Longrightarrow}_{\overline{\mu},\overline{\mu}^{-1}\in L^{\infty}_{\mathrm{loc}}}\overline{\mu}\in W^{1,\infty}_{\mathrm{loc}}(\mathbb{R}^{d}).$
The following result will play a key role in obtaining a proper bound for the viscosity term.
Lemma 4.3.
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu\in W^{1,\infty}_{\mathrm{loc}}(\mathbb{R}^{d})$ and $\mu^{-1}\in L^{\infty}_{\text{loc}}(\mathbb{R}^{d})$. Fix $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
(4.2) $\int_{\mathbb{R}^{d}}\big<\nabla\log\mu(x),x-T^{\nu}_{\mu}(x)\big>d\mu(x)\leq d.$
Proof.
From the regularity assumptions on $\mu$, $\text{supp}(\mu)=\mathbb{R}^{d}.$. Therefore, it is sufficient to prove
$\int_{\mathbb{R}^{d}}\big<\nabla\mu(x),x-T_{\mu}^{\nu}(x)\big>dx\leq d.$
By Brenier Theorem, see [Bre, Proposition  3.1]) and the fact that $\text{supp}(\mu)=\mathbb{R}^{d}$, we have that $T_{\mu}^{\nu}=\nabla\phi$ $\mu$-a.e, where $\phi:\mathbb{R}^{d}\to\mathbb{R}$ is a finite convex function. Being $\phi$ locally Lipschitz, $T^{\nu}_{\mu}$ is locally bounded. We then regularize the transport map by a convolution kernel $\rho_{\alpha}:=\alpha^{-d}\rho(\frac{x}{\alpha})$, where $\rho\in C^{\infty}(\mathbb{R}^{d})$ is a probability density, compactly supported in the unitary ball $B_{1}(0)$ and $\alpha>0$. Set $T_{\alpha}:=\nabla\phi*\rho_{\alpha}=\nabla(\phi*\rho_{\alpha})$, then $T_{\alpha}$ inherits the local boundedness of $T_{\mu}^{\nu}$, and $T_{\alpha}$ converges to $T_{\mu}^{\nu}$ in $L^{2}_{\text{loc}}(\mathbb{R}^{d})$ (see [Evans, Theorem  6 pag. 630]). Since convolution preserves convexity, $\phi_{\alpha}:=\phi*\rho_{\alpha}$ is convex. Therefore, the following inequality holds for every $\alpha>0$
$\operatorname{div}(T_{\alpha}(x))=\text{Trace}(D^{2}\phi_{\alpha}(x))\geq 0\quad\forall x\in\mathbb{R}^{d}.$
We can then integrate this inequality against non-negative, compactly supported Lipschitz functions to obtain, for all $\alpha>0$,
[…]
Since test functions are compactly supported, we can pass to the limit as $\alpha\to 0$ to obtain
[…]
Consider now for $R>0$, $\chi_{R}(x):=\chi(\frac{x}{R})$, where $\chi\in C^{\infty}_{c}(\mathbb{R}^{d})$ is constantly equal to 1 on $B_{1}(0)$, $\chi\leq 1$ on $B_{2}(0)\setminus B_{1}(0)$ and $\chi$ equal to zero on $\mathbb{R}^{d}\setminus B_{2}(0)$. We test inequality (4.3) with the non negative function $\xi=\chi_{R}\mu\in W^{1,\infty}_{c}(\mathbb{R}^{d})$ and we obtain
$-\int_{\mathbb{R}^{d}}\big<x-T_{\mu}^{\nu}(x),\nabla\chi_{R}(x)\big>d\mu(x)$ $-\int_{\mathbb{R}^{d}}\big<x-T_{\mu}^{\nu}(x),\nabla\mu(x)\big>\chi_{R}(x)dx=$
$=-\int_{\mathbb{R}^{d}}\big<x-T_{\mu}^{\nu}(x),\nabla\xi(x)\big>dx$
$=\int_{\mathbb{R}^{d}}\text{div}(x)\xi(x)dx+\int_{\mathbb{R}^{d}}\big<T_{\mu}^{\nu}(x),\nabla\xi(x)\big>dx$
[…]
On the other hand, we have the following bound
$\int_{\mathbb{R}^{d}}\big<x-T_{\mu}^{\nu}(x),\nabla\chi_{R}(x)\big>d\mu(x)\leq\frac{C}{R}\Big(\int_{\mathbb{R}^{d}}|x-T_{\mu}^{\nu}(x)|^{2}d\mu(x)\Big)^{\frac{1}{2}}=\frac{C}{R}W_{2}(\mu,\nu)$
that combined with the previous inequality gives
$\int_{\mathbb{R}^{d}}\big<x-T_{\mu}^{\nu}(x),\nabla\mu(x)\big>\chi_{R}(x)dx\leq\int_{\mathbb{R}^{d}}\big<x-T_{\mu}^{\nu}(x),\nabla\chi_{R}(x)\big>d\mu(x)+d\leq\frac{C}{R}W_{2}(\mu,\nu)+d.$
Passing to the limit $R\to\infty$, we get the claim.
Remark 4.4.
The inequality (4.2) has a Riemannian flavour: In a $d$-dimensional smooth Riemannian manifold $(M,g)$ with non-negative Ricci curvature we have
$\Delta_{g}\frac{1}{2}d^{2}(x,y)\leq d,$
where $d(x,y)$ is the geodesic distance between $x$ and $y$ in $M$. See [petersen2006riemannian, Lemma  42]. In our context, the above inequality can be interpret as a weak analogue of
$\Delta_{\mu}\frac{1}{2}W_{2}^{2}(\mu,\nu):=\int_{\mathbb{R}^{d}}\Delta(\mathcal{D}_{\mu}\frac{1}{2}W_{2}^{2}(\mu,\nu)))d\mu\leq d.$
The inequality stated in Lemma 4.3 is a particular case of a more general result of independent interest, at least in the case $d\geq 2$.
Proposition 4.5.
Suppose $d\geq 2$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu^{-1}\in L^{\infty}_{\text{loc}}$ and $\mu\in W^{1,\infty}_{\text{loc}}$. Let $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}$ be a $\Lambda$-geodesically concave functional and suppose $\mathcal{F}$ to be flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
$\int_{\mathbb{R}^{d}}\big<\nabla\log\mu,\nabla\mathcal{D}_{\mu}\mathcal{F}(\mu,x)\big>d\mu\leq d\Lambda.$
Recall that the squared Wasserstein distance is $1$-geodesically concave, i.e. the map $\mu\mapsto W_{2}^{2}(\mu,\nu)$ is 1-concave along Wasserstein geodesics for every $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Equivalently, $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a geodesic PC space, see [AGS, Definition 12.3.1] and later discussion. When $\mu$ is absolutely continuous, its flat derivative is given (up to an additive constant) by the optimal Kantorovich potential $\phi$ associated with the transport from $\mu$ to $\nu$. In particular, denoting by $T_{\mu}^{\nu}=\nabla\phi$ the optimal transport map, we have
$\mathcal{D}_{\mu}W_{2}^{2}(\cdot,\nu)(\mu,x)=\frac{|x|^{2}}{2}-\phi(x),$
that explains why this proposition is a generalization in dimension $d\geq 2$ of Lemma 4.3.
Firstly, we derive a sufficient condition for a function on the Wasserstein space to have $\Lambda$-concave flat derivative, namely
Lemma 4.6.
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a flat differentiable function at a $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, s.t. there exists $\Lambda=\Lambda(\mu)$ for which the following inequality holds
(4.4) $\mathcal{F}((1-h)\mu+h\delta_{y_{t}})\leq(1-t)\mathcal{F}((1-h)\mu+h\delta_{y_{0}})+t\mathcal{F}((1-h)\mu+h\delta_{y_{1}})-h\Lambda\frac{t(1-t)}{2}|y_{0}-y_{1}|^{2},$
for all $y_{0},y_{1}\in\mathrm{supp}(\mu)$, with $y_{t}:=(1-t)y_{0}+ty_{1}\in\mathrm{supp}(\mu),\,t,h\in[0,1].$
Then $x\mapsto D_{\mu}\mathcal{F}(\mu,x)$ is $\Lambda$-convex in the classical $\mathbb{R}^{d}$ sense on $\mathrm{supp}(\mu)$.
Proof.
We derive the desired property from the explicit formula given in (2.12) for the flat derivative. Indeed, consider $y_{0},y_{1}\in\text{supp}(\mu)$ s.t. their convex interpolation $y_{t}=(1-t)y_{0}+ty_{1}\in\text{supp}(\mu)$ for all $t\in[0,1]$. Thanks to the assumption (4.4) on $\mathcal{F}$, and the representation formula (2.12), we have
$\mathcal{D}_{\mu}\mathcal{F}(\mu,y_{t})$ $=\lim_{h\to 0}\frac{\mathcal{F}((1-h)\mu+h\delta_{y_{t}})-\mathcal{F}(\mu)}{h}$
$\leq(1-t)\lim_{h\to 0}\frac{\mathcal{F}((1-h)\mu+h\delta_{y_{0}})-\mathcal{F}(\mu)}{h}+t\lim_{h\to 0}\frac{\mathcal{F}((1-h)\mu+h\delta_{y_{1}})-\mathcal{F}(\mu)}{h}$
$-\lim_{h\to 0}\frac{h{\frac{\Lambda}{2}t(1-t)|y_{0}-y_{1}|^{2}}}{h}$
$=(1-t)\mathcal{D}_{\mu}\mathcal{F}(\mu,y_{0})+t\mathcal{D}_{\mu}\mathcal{F}(\mu,y_{1})-{\frac{\Lambda}{2}t(1-t)|y_{0}-y_{1}|^{2}}.$
We now consider $\Lambda$-geodesically convex functions in the sense of the following definition.
Definition 4.7.
Given $\Lambda\in\mathbb{R}$, we say that $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is $\Lambda$-geodesically convex w.r.t. $W_{2}$, if for any $\mu_{0},\mu_{1}\in\text{Dom}(\mathcal{F})$ and for every $W_{2}$ constant speed geodesic $t\mapsto\mu_{t}\in\text{Dom}(\mathcal{F})$ connecting $\mu_{0}$ and $\mu_{1}$, $\mathcal{F}$ is $\Lambda$-convex along this geodesic, i.e.
$\mathcal{F}(\mu_{t})\leq(1-t)\mathcal{F}(\mu_{0})+t\mathcal{F}(\mu_{1})-\frac{\Lambda}{2}t(1-t)W_{2}^{2}(\mu_{0},\mu_{1}),\quad\forall t\in[0,1].$
We say that $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{-\infty\}$ is $\Lambda$-geodesically concave w.r.t. $W_{2}$ if F-F is $(-\Lambda)$-geodesically convex.
The following Lemma shows a geodesic convexity property along mixtures, i.e. flat convex combination of measures.
Lemma 4.8.
Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continuous $\Lambda$-geodesically convex function. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then
  • $\forall h\in[0,1]$ the function
    $\nu\mapsto\mathcal{F}((1-h)\mu+h\nu)$
    is $h\Lambda$-geodesically convex.
  • In addition, if $\mathcal{F}$ is flat differentiable at $\mu$, then $\mathcal{F}$ satisfies the assumption (4.4) of Lemma 4.6.
Proof.
The proof relies on the result contained in [CavSavSod23, Proposition  9.1] that states: In dimension $d\geq 2$, if $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a continuous $\Lambda$-geodesically convex function, then it is totally geodesically convex, namely given $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and any $\gamma\in\Gamma(\mu_{0},\mu_{1})$ we have
(4.5) $\mathcal{F}(\mu_{t})\leq(1-t)\mathcal{F}(\mu_{0})+\mathcal{F}(\mu_{1})-\frac{\Lambda}{2}t(1-t)(W_{2}^{\gamma}(\mu_{0},\mu_{1}))^{2}\quad\forall t\in[0,1],$
where $\mu_{t}=(\pi_{1}+t(\pi_{2}-\pi_{1}))_{\#}\gamma$, and $W_{2}^{\gamma}(\mu_{0},\mu_{1})^{2}=\int_{\mathbb{R}^{2d}}|y-x|^{2}d\gamma(x,y)$ is defined in (2.2).
Now, given $\nu_{0},\nu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we consider the mixture $\tau_{h,t}:=(1-h)\mu+h\nu_{t}$, defined for $h\in[0,1]$ and $t\in[0,1]$, where $\nu_{t}:=(\pi_{1}+t(\pi_{2}-\pi_{1})_{\#}\widetilde{\gamma}$ is the constant speed geodesic induced by the optimal plan $\widetilde{\gamma}\in\Gamma_{0}(\nu_{0},\nu_{1})$, and choose the plan
$\gamma_{h}:=(1-h)(\text{Id}\times\text{Id})_{\#}\mu+h\widetilde{\gamma}$.
Observe that $\gamma_{h}\in\Gamma(\tau_{h,0},\tau_{h,1})$, and $\tau_{h,t}=(\pi_{1}+t(\pi_{2}-\pi_{1}))_{\#}\gamma_{h}$.
Therefore, the total convexity property $\mathcal{F}$ yields
$\mathcal{F}(\tau_{h,t})\leq(1-t)\mathcal{F}(\tau_{h,0})+t\mathcal{F}(\tau_{h,1})-\frac{\Lambda}{2}t(1-t)(W^{\widetilde{\gamma}}_{2}(\tau_{h,0},\tau_{h,1}))^{2}.$
then (1) follows observing that
$(W^{\widetilde{\gamma}}_{2}(\tau_{h,0},\tau_{h,1}))^{2}=hW^{2}_{2}(\nu_{0},\nu_{1}).$
Assertion (2) follows by choosing for all $y_{0},y_{1}\in\mathrm{supp}(\mu)$ the constant speed geodesic $\nu_{t}=\delta_{y_{t}}$, whenever $y_{t}\in\text{supp}(\mu)$, connecting $\nu_{0}=\delta_{y_{0}}$ and $\nu_{1}=\delta_{y_{1}}$: indeed in this case $\nu_{t}=\delta_{y_{t}}$ where $y_{t}=(1-t)y_{0}+ty_{1}$, and the $h\Lambda$-geodesically convexity of $\nu\mapsto\mathcal{F}((1-h)\mu+h\nu)$ restricted to Dirac masses corresponds to (4.4).
Remark 4.9.
We are not able to establish the previous result in dimension $d=1$, since our argument strongly relies on [CavSavSod23, Proposition 9.1], which is only available for $d>1$.
The main difficulty lies in the fact that, in general, even for $d>1$, the curve
$t\in[0,1]\longmapsto\tau_{h,t}=(1-h)\mu+h\nu_{t}$
is not a $W_{2}$-geodesic, even if $t\mapsto\nu_{t}$ is one. The total convexity property avoid this issue.
Corollary 4.10.
Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continous $\Lambda$-geodesically convex function. If $\mathcal{F}$ is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ then $x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is convex in $\mathrm{supp}(\mu)$.
We are now in the position to prove the announced Proposition 4.5.
Proof of Proposition 4.5.
The argument of the proof follows the same lines of the proof of Lemma 4.3. In that context, the 1-concave function $\frac{x^{2}}{2}-\phi(x)=\mathcal{D}_{\mu}\frac{W^{2}(\mu,\nu)(x)}{2}$ is now replaced by the $\Lambda$-concave function $\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$. ∎
We conclude this section with an application of the above results.
Lemma 4.11.
Let $U^{\varepsilon}$ be a classical subsolution of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}(x)$. If $U^{\varepsilon}$ is s.t. $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)\leq C$ for some $C$ independent of $(t,\mu,x)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$, then $U^{\varepsilon}-\varepsilon C(t-T)$ is a classical subsolution of (1.1). In particular, provided that $d\geq 2$, $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)\leq C$ is satisfied by a $\Lambda$-geodesically concave functional, where $C=d\Lambda$.
Proof.
If $U^{\varepsilon}$ is a classical subsolution of (1.2), then $U^{\varepsilon}$ is $C^{2}$. In particular, $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)$ is well defined in $\text{supp}(\mu)$. Moreover,
$-\partial_{t}U^{\varepsilon}+\mathcal{H}((\mathrm{Id}\times\partial_{\mu}U^{\varepsilon})_{\#}\mu)$ $=-\partial_{t}U^{\varepsilon}+\int_{\mathbb{R}^{d}}H(x,\mu,\partial_{\mu}U^{\varepsilon})d\mu$
$\underbrace{\leq}_{U^{\varepsilon}\text{classical subsolution}}\varepsilon\int_{\mathbb{R}^{d}}\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)d\mu\leq\varepsilon C.$
The second part of the statement is an immediate consequence of Corollary 4.10. ∎
5. Vanishing Viscosity Limit
In this section, we state and prove the vanishing viscosity result.
Theorem 5.1.
Let $H\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ and $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ satisfy the assumption 2.15.
Let $U^{\varepsilon}$ and $U$ be two bounded viscosity solutions of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}(x)$, and (1.1), respectively. Assume that $U^{\varepsilon}(T,\mu)=U(T,\mu)=\mathcal{G}(\mu)$. Additionally, suppose $U^{\varepsilon}$ to be $W_{1}$-Lipschitz continuous, and $U$ to be $W_{2}$-Lipschitz continuous, Lipschitz in time, and continuous w.r.t. the weak convergence. Then there exists a constant $C\geq 0$ depending only on the data s.t.
$\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}|U^{\varepsilon}(t,\mu)-U(t,\mu)|\leq C\sqrt{\varepsilon}.$
Proof.
The first part of the proof is mainly inspired by the proof of [Alekos, Theorem  2.10].
Firstly, we focus on the upper bound: it will be clear through the proof that a symmetry argument, inverting the role of the two functions, allows us to concentrate on one of the two sides of convergence.
Step 1: Error quantification term.
Let
(5.1) $E^{+}_{\varepsilon}:=\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}U^{\varepsilon}(t,\mu)-U(t,\mu)$
be the one side error between the two solutions. We want to prove that
$E^{+}_{\varepsilon}\leq C\sqrt{\varepsilon}.$
We note that $E^{+}_{\varepsilon}$ is bounded by the boundedness assumptions on the functions. Clearly, if $E^{+}_{\varepsilon}\leq 0$ then (5.1) holds. Therefore, we assume $E^{+}_{\varepsilon}>0$.
Step 2: Doubling variables penalization.
We invoke the usual doubling variables arguments by considering the penalized function
$\Phi_{\delta,\alpha}(s,t,\mu,\nu)=U^{\varepsilon}(s,\mu)-\delta(\mathcal{E}(\mu)+\pi\mathcal{M}_{2}(\mu))-U(t,\nu)-\delta\pi\mathcal{M}_{2}(\nu)-\frac{1}{2\alpha}W_{2}^{2}(\mu,\nu)-\frac{1}{2\alpha}|t-s|^{2}-\frac{2T-t-s}{4T}E^{+}_{\varepsilon}$,
over $0,T]\times[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$. Here $\alpha,\delta>0$, with $\alpha$ to be chosen in terms of $\varepsilon$.
Now, being $\mathcal{E}^{*}:=\mathcal{E}+\pi\mathcal{M}_{2}\geq 0$, we have that the function
$\delta\mapsto M_{\delta,\alpha,\eta}:=\max_{(s,t,\mu,\nu)\in[0,T]^{2}\times(\mathcal{P}_{2}(\mathbb{R}^{d}))^{2}}\Phi_{\delta,\alpha,\eta}(s,t,\mu,\nu)$
is nonincreasing. Therefore, for fixed $\alpha>0$, the limit
$M_{\alpha}:=\lim_{\delta\to 0}M_{\delta,\alpha}$
exists. Moreover, from the inequality
$M_{\delta,\alpha,\eta}=\Phi_{\delta,\alpha,\eta}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})\leq M_{\delta/2,\alpha,\eta}-\frac{\delta}{2}(\mathcal{E}(\overline{\mu})+\pi\mathcal{M}_{2}(\overline{\mu}))-\frac{\delta\pi}{2}\mathcal{M}_{2}(\overline{\nu}).$
we deduce
$\frac{\delta}{2}(\mathcal{E}(\overline{\mu})+\pi\mathcal{M}_{2}(\overline{\mu}))+\frac{\delta\pi}{2}\mathcal{M}_{2}(\overline{\nu})\leq M_{\delta/2,\alpha,\eta}-M_{\delta,\alpha,\eta}\xrightarrow{\delta\to 0}0$
for fixed $\alpha,\eta>0$. In particular,
$\lim_{\delta\to 0}\delta\mathcal{E}^{*}(\overline{\mu})=0.$
Moreover, since $\mathcal{E}^{*}$ satisfies $\mathcal{E}^{*}(\overline{\mu})\geq\frac{\pi}{2}\mathcal{M}_{2}(\overline{\mu})-\frac{d}{2}\log 2$, for fixed $\alpha>0$,
(5.2) $\lim_{\delta\to 0}\delta(\mathcal{M}_{2}(\overline{\mu})+\mathcal{M}_{2}(\overline{\nu}))=0.$
Therefore, both entropy and moment penalizations vanish with $\delta$.
Now, rearranging the terms in the inequality $\Phi_{\delta,\alpha,\eta}(\overline{s},\overline{t},\overline{\mu},\overline{\mu})\leq\Phi_{\delta,\alpha,\eta}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})$, leads to the inequality
$\frac{1}{2\alpha}W^{2}_{2}(\overline{\mu},\overline{\nu})$ $\leq U^{\varepsilon}(\overline{s},\overline{\mu})-U^{\varepsilon}(\overline{s},\overline{\mu})+U(\overline{t},\overline{\mu})-U(\overline{t},\overline{\nu})$
$=U(\overline{t},\overline{\mu})-U(\overline{t},\overline{\nu})$
By the $W_{2}$-Lipschitz regularity of $U$, we recover the following bound
$\frac{1}{2\alpha}W^{2}_{2}(\overline{\mu},\overline{\nu})\leq\mathrm{Lip}(U;W_{2})W_{2}(\overline{\mu},\overline{\nu}).$
In particular,
(5.3) $W_{2}(\overline{\mu},\overline{\nu})\leq 2\alpha\mathrm{Lip}(U;W_{2}).$
We have a similar estimate in time as a consequence of the following inequality $\Phi_{\delta,\alpha}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})\leq\Phi_{\delta,\alpha}(\overline{s},\overline{s},\overline{\mu},\overline{\nu})$ that yields
$\frac{|\overline{t}-\overline{s}|^{2}}{2\alpha}\leq U(\overline{t},\overline{\nu})-U(\overline{s},\overline{\nu})+\frac{\overline{t}-\overline{s}}{4T}E_{\varepsilon}^{+},$
which implies, thanks to the Lipschitz continuity in time of $U$ and the boundedness of $E_{\varepsilon}^{+}$,
(5.4) $|\overline{t}-\overline{s}|\leq C\alpha.$
Step 3: Error estimate at the terminal time.
We now consider the three cases: either $\overline{s}=T$, or $\overline{t}=T$, or $\overline{t},\overline{s}<T.$
The case $(\overline{s},\overline{t})\in\{T\}\times[0,T].$ We consider the inequality $\Phi_{\delta,\alpha}(t,t,\mu,\mu)\leq\Phi_{\delta,\alpha}(\overline{s},\overline{t},\overline{\mu},,\overline{\nu})$ $\forall(t,\mu)\in[0,T]\times\mathrm{Dom}(\mathcal{E})$, we get
$U^{\varepsilon}(t,\mu)-U(t,\mu)-\delta\mathcal{E}(\mu)-2\delta\pi\mathcal{M}(\mu)$ $\leq\mathcal{G}(\overline{\mu})-U(\overline{t},\overline{\nu})+\frac{2T-2t}{4T}E_{\varepsilon}^{+}$
$\leq\mathcal{G}(\overline{\mu})-\mathcal{G}(\overline{\nu})+\mathcal{G}(\overline{\nu})-U(\overline{t},\overline{\nu})+\frac{2T-2t}{4T}E_{\varepsilon}^{+}$
$\leq\mathrm{Lip}(\mathcal{G};W_{2})W_{2}(\overline{\mu},\overline{\nu})+\mathrm{Lip}_{t}(U)|T-\overline{t}|+\frac{2T-2t}{4T}E_{\varepsilon}^{+}$
[…]
for all $(t,\mu)\in[0,T]\times\mathrm{Dom}(\mathcal{E})$. Having obtained a bound from above that does not depend on $\delta$ we can send it to zero. Therefore,
$U^{\varepsilon}(t,\mu)-U(t,\mu)\leq C\alpha+\frac{1}{2}E_{\varepsilon}^{+}\quad\forall(t,\mu)\in[0,T]\times\mathrm{Dom}(\mathcal{E}).$
Passing to the supremum in the LHS and recalling that $\overline{\mathrm{Dom}(\mathcal{E})}^{W_{2}}=\mathcal{P}_{2}(\mathbb{R}^{d})$ we get
$E_{\varepsilon}^{+}\leq C\alpha+\frac{1}{2}E_{\varepsilon}^{+}.$
Finally, the choice $\alpha=\sqrt{\varepsilon}$ gives the claim in this scenario.
The case $(\overline{s},\overline{t})\in[0,T]\times\{T\}$ is completely analogous.
Step 4: The case $(\overline{s},\overline{t})\in[0,T)\times[0,T)$: Application of the definitions of viscosity solution.
By adapting Lemma 4.2 to functions dependent on time,
$\begin{cases}&(\frac{\overline{s}-\overline{t}}{\alpha}-\frac{1}{4T}E_{\varepsilon}^{+},\mu\otimes\delta_{\frac{1}{\alpha}(x-T_{\overline{\mu}}^{\overline{\nu}}(x))+2\pi\delta x})\in\partial^{+}(U^{\varepsilon}-\delta\mathcal{E})(\overline{s},\overline{\mu})\\ &(\frac{\overline{s}-\overline{t}}{\alpha}+\frac{1}{4T}E_{\varepsilon}^{+},\overline{\Sigma})\in\partial^{-}U(\overline{t},\overline{\nu}),\end{cases}$
where $\overline{\Sigma}=(\pi_{1},\frac{\pi_{1}-\pi_{2}}{\alpha}-2\pi\delta\pi_{2})_{\#}\overline{\sigma}$, with $\overline{\sigma}\in\Gamma_{0}(\overline{\nu},\overline{\mu})$ s.t. $(\pi_{2},\pi_{1})_{\#}\overline{\sigma}=(\mathrm{Id}\times T_{\overline{\mu}}^{\overline{\nu}})_{\#}\overline{\mu}$. Moreover, we have $\mathcal{I}(\overline{\mu})<\infty$. Therefore, we can apply the Definitions 2.18 and 2.16 to obtain the following inequalities
$\frac{1}{4T}E_{\varepsilon}^{+}-\frac{\overline{s}-\overline{t}}{\eta}$ $+\int_{\mathbb{R}^{d}}H\left(x,\left(\frac{1}{\alpha}(x-T_{\overline{\mu}}^{\overline{\nu}}(x))+2\pi\delta x\right),\overline{\mu}\right)d\overline{\mu}(x)$
$+\varepsilon\int_{\mathbb{R}^{d}}\nabla\log\overline{\mu}(x)\cdot\left(\frac{1}{\alpha}\left(x-T_{\overline{\mu}}^{\overline{\nu}}(x)\right)+2\pi\delta x\right)d\overline{\mu}(x)\leq K(\varepsilon)\delta,$
and
$-\frac{1}{4T}E_{\varepsilon}^{+}{-}\frac{\overline{s}-\overline{t}}{\eta}$ $+\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}H\left(y,\frac{1}{\alpha}(x-y)-2\pi\delta y,\overline{\nu}\right)d\overline{\sigma}(y,x)\geq 0.$
Here, we call $K(\varepsilon)$ the penalization constant in the Definition 2.18, stressing its possible dependence on $\varepsilon$. Subtracting the two inequalities, we obtain the following.
$\frac{1}{2T}E_{\varepsilon}^{+}\leq K(\varepsilon)\delta+I_{1}+I_{2}+I_{3},$
where
$I_{1}:=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left[H\left(y,\frac{1}{\alpha}(x-y)+2\pi\delta y,\overline{\nu}\right)-H\left(x,\left(\frac{1}{\alpha}(x-y)+2\pi\delta x\right),\overline{\mu}\right)\right]d\overline{\sigma}(y,x),$
$I_{2}$ $:=-\varepsilon\int_{\mathbb{R}^{d}}\nabla\log\overline{\mu}(x)\cdot\frac{1}{\alpha}(x-T_{\overline{\mu}}^{\overline{\nu}}(x))d\overline{\mu}(x),$
and
$I_{3}:=-2\pi\delta\varepsilon\int_{\mathbb{R}^{d}}\nabla\log\overline{\mu}(x)\cdot xd\overline{\mu}(x).$
In what follows, the constant $c$ denotes a constant that changes from line to line, and is independent of $\alpha,\delta$ and $\varepsilon$.
Step 5: Estimate of the integrals $I_{1}\&I_{3}.$
Using Cauchy-Schwarz inequality and (5.2), (5.3) we get
$|I_{1}|$ $\leq c\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\Big(1+\left|\frac{x-y}{\alpha}\right|+\delta|x|+\delta|y|\Big)\Big(|x-y|+\delta|y|+\delta|x|+W_{2}(\overline{\mu},\overline{\nu})\Big)d\overline{\sigma}$
$\leq cW_{2}(\overline{\mu},\overline{\nu})+\frac{W^{2}_{2}(\overline{\mu},\overline{\nu})}{2\alpha}+\delta(\mathcal{M}^{\frac{1}{2}}(\overline{\mu})+\mathcal{M}^{\frac{1}{2}}(\overline{\nu}))+\delta(\mathcal{M}(\overline{\mu})+\mathcal{M}(\overline{\nu}))$
[…]
The third integral $I_{3}$ can be bound thanks to the fact that $\overline{\mu}\in W_{\mathrm{loc}}^{1,\infty}$and $\overline{\mu}^{-1}\in L_{\mathrm{loc}}^{\infty}$. Indeed, in this case, as in the proof in [DS, Theorem  2.14], we can perform an integration by parts and prove
$I_{3}=2\pi\delta\varepsilon\int_{\mathbb{R}^{d}}\operatorname{div}(x)d\overline{\mu}(x)=2\pi d\delta\varepsilon\leq c\delta\varepsilon.$
Step 6: Estimate of the viscosity term.
Adapting Lemma 4.2 to functions depending on time, the measure $\overline{\mu}$ has enough regularity to invoke Lemma 4.3. Therefore, we obtain
$I_{2}=-\varepsilon\int_{\mathbb{R}^{d}}\big<\nabla\log\overline{\mu},\frac{x-T^{\overline{\nu}}_{\overline{\mu}}(x)}{\alpha}\big>d\overline{\mu}\leq\frac{\varepsilon}{\alpha}d.$
Step 7: Conclusion.
In summary, the previous estimates give
$\frac{1}{2T}E_{\varepsilon}^{+}\leq K(\varepsilon)\delta+c(\alpha+\frac{\varepsilon}{\alpha})+\delta^{\frac{1}{2}}+\delta(\mathcal{M}(\overline{\mu})+\mathcal{M}(\overline{\nu})),$
where the constant $c$ is independent of $\varepsilon,\alpha,\delta$. Now we pass to the limit $\delta\to 0$ and, recalling the vanishing behavior of the penalizations (5.2), we get
$\frac{1}{2T}E_{\varepsilon}^{+}\leq c(\alpha+\frac{\varepsilon}{\alpha})$
We then optimize w.r.t. $\alpha$, so that $\alpha=\sqrt{\varepsilon}$. With this choice of the we finally get
$\frac{1}{2T}E_{\varepsilon}^{+}\leq c\sqrt{\varepsilon}.$
By the considerations at the beginning of the proof, the same argument allows to obtain the bound
$E_{\varepsilon}^{-}:=\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}U(t,\mu)-U^{\varepsilon}(t,\mu)\leq C\sqrt{\varepsilon}.$
Remark 5.2.
The boundedness assumption on the functions can be relaxed to allow for quadratic growth, as already observed in [DS]. Indeed, in the proof it suffices to introduce an additional moment penalization to compensate for the growth of the difference $U^{\varepsilon}-U$.
Remark 5.3.
As in the finite-dimensional setting, the rate of convergence follows from the regularity assumptions on the Hamiltonian: modifying its local modulus of continuity directly affects the convergence speed. The argument can be adapted to this framework without additional difficulty. Finally, we can also handle the case of terminal costs depending on the viscosity parameter $\varepsilon$ thanks to the stability property established in [DS, Theorem  4.1].
Remark 5.4.
The above method yields an optimal convergence rate. For a Hamiltonian satisfying 2.15, this rate is the best one can expect without assuming convexity of $H$, since the finite-dimensional setting is embedded in this framework. It would be interesting to investigate whether this bound can be improved in the convex case, as independently shown in [cirant2025convergenceratesvanishingviscosity, chaintron2025optimalrateconvergencevanishing].
We end this section with a one-sided improvement of the previous bound in a more restrictive case.
Proposition 5.5.
Under the assumption of Theorem 5.1, if $U^{\varepsilon}$ is a classical solution of (1.2) s.t. $\Delta\mathcal{D}_{\mu}U^{\varepsilon}\leq C$ where $C:[0,T]\to[0,+\infty)$ Then
$U^{\varepsilon}(t,\mu)-U(t,\mu)\leq C(T-t)\varepsilon.$
In particular, if $d\geq 2$, and $\mu\mapsto U^{\varepsilon}(t,\mu)$ is geodesically $\Lambda$-concave, the result holds with $C=d\Lambda.$
Proof.
It is sufficient to observe that we can apply Lemma 4.11 and then the comparison principle in [BSOT, Theorem  2.11]. ∎
6. Acknowledgments
This research is supported by the project King Abdullah University of Science and Technology (KAUST) ORA-CRG2021-4674 “Mean-Field Games: models, theory and computational aspects”; and by the project PRIN 2022 (prot. 2022W58BJ5) “PDEs and optimal control methods in mean field games, population dynamics and multi-agent models”. While this work was written GCS and DT were associated to INdAM and the group GNAMPA. The authors thank Charles Bertucci and Alekos Cecchin for the fruitful discussions.